### v2

# Achieving Optimal Sample Complexity for a Broader Class of Signals in Sparse Phase Retrieval

Abstract
Sparse phase retrieval, which aims to recover a $k$-sparse signal from $m$ phaseless measurements, poses a fundamental question regarding the minimal sample complexity required for success. While the optimal sample complexity is known to be $\Omega(k\log n)$, existing algorithms can only achieve it for signals under restrictive structural assumptions. This paper introduces a novel and robust initialization algorithm, termed generalized Exponential Spectral Pursuit (gESP), designed to overcome this limitation. Theoretically, we prove that gESP significantly expands the family of signals that can be recovered with the optimal sample complexity. Our analysis unifies existing results on previously studied signal models and surpasses them by establishing performance bounds for a more general class of signals. Extensive simulations validate our theoretical findings, demonstrating that gESP consistently outperforms state-of-the-art methods across diverse signal types, thereby pushing the boundaries of efficient and optimal sparse phase retrieval.
Index Terms:
Sparse phase retrieval, sample complexity, sparsity, information-theoretical bound.
I Introduction
In many applications, only the intensity information is retained by the measurement system due to hardware limitations [1]. For example, in optical imaging [2, 3] and X-ray crystallography [4, 5], common sensors such as charge-coupled devices, photosensitive films and human eyes merely record the amplitude of the light wave but ignore the phase. In these settings, one is faced with the task of recovering the original signal from its magnitude-only measurements, which is often referred to as phase retrieval (PR) [6, 7, 8, 9, 10]. Succinctly stated, the main task of PR is to reconstruct the $n$-dimensional signal $\mathbf{x}\in\mathbb{C}^{n}$ from a group of phaseless measurements $\mathbf{y}\in\mathbb{R}^{m}$ with
$y_{i}=|\mathbf{a}_{i}^{*}\mathbf{x}|,\ i=1,2,\cdots m,$ (1)
where $\mathbf{a}_{i}\in\mathbb{C}^{n}$ is the measurement vector. In this paper, we consider the complex Gaussian measurements, i.e., $\mathbf{a}_{i}\sim\mathcal{CN}(n)$, and $\mathbf{x}$ is $k$-sparse, which means that there are at most $k$ non-zero entries in $\mathbf{x}$. This problem is the so-called sparse phase retrieval problem. Researchers are interested in the minimum measurement number that one needs to reconstruct the signal [11, 12, 13, 14]. It has been shown that with m=4k2m=4k-2 phaseless measurements [15], the non-linear system (1) with a $k$-sparse prior produces the unique solution based on generic sensing measurements. Moreover, researches have proved that
$m=\Omega(k\log n)$ (2)
complex Gaussian measurements are sufficient to guarantee the stability of the solution [16, 17]. This sample complexity is often called the information-theoretical bound for sparse phase retrieval.
For polynomial algorithms, they appear to require more measurements for the exact recovery. Back in 2013, Li and Voroninski [18] utilized the matrix lifting technique and convex relaxation methods to solve the sparse phase retrieval problem. They proved that it requires
$m=\Omega(k^{2}\log n)$ (3)
measurements to produce an exact reconstruction. Moreover, they further proved that this requirement cannot be reduced to be smaller than $m=\Omega({k^{2}}/{\log^{2}n})$ for their supposed model. While their sample complexity revealed a clear discrepancy compared to (2), the authors were unsure whether this difference was inherent to the problem itself or simply a consequence of limitations in their modeling approach. This is afterwards called the statistical-to-computational gap.
This gap is further investigated in the thresholded Wirtinger Flow (TWF) algorithm [19], which is a typical greedy algorithm containing two stages — the initialization stage and the refinement stage. They showed that both stages require $m=\Omega(k^{2}\log n)$ samples for the exact final recovery. This two-stage approach was followed in subsequent algorithms, e.g., SPARse Truncated Amplitude flow (SPARTA[20], Compressive Phase Retrieval with Alternating Minimization (CoPRAM[21], and Subspace Phase Retrieval (SPR[22]. It is shown via some geometric analysis [23, 22, 24] and sophisticated optimization approaches [25] that the sample complexity of the refinement stage can be improved to the information-theoretical bound (2), if a $\delta$-neighborhood estimate of the target signal is provided. However, to achieve this, the initialization stage requires
$m=\Omega(k^{2}\log n)$ (4)
measurements, which is still far from the information-theoretical bound (2). It has become a consensus that the statistical-to-computational gap results from the initialization stage [23, 26]. More precisely, it is the requirement of finding an estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ that leads to the statistical-to-computational gap.
Despite the persistence of this gap, numerous efforts are underway to achieve improved results. For the sake of notational simplicity, we define $s(p)$ as
$s(p)=\frac{\|\mathbf{x}\|^{2}}{\sum_{j=1}^{p}|x_{(j)}|^{2}},~~p=1,2,\cdots,n,$ (5)
where $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ are the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. A formal definition and the corresponding interpretation are presented in Section III-A. Wu and Rebeschini [27] presented the algorithm call Hadamard Wirtingle Flow (HWF). They showed that the gap vanished for some specific structure of the target signal $\mathbf{x}$. Specifically, by assuming that the minimum magnitude of non-zeros $x_{(k)}$ is on the order of $\frac{\|\mathbf{x}\|}{\sqrt{k}}$, i.e.,
$\left\lvert x_{(k)}\right\rvert=\Omega\left(\frac{\|\mathbf{x}\|}{\sqrt{k}}\right),$ (6)
they show that the sample complexity can be reduced to
$m=\Omega\left(\max\left\{ks(1)\log n,\sqrt{ks(1)}\log^{3}n\right\}\right).$ (7)
This means that when the maximum entry $x_{(1)}$ of $\mathbf{x}$ is on the order of $\|\mathbf{x}\|$ (ignoring the $\log^{3}n$ term), (7) can be further reduced to (2), and thus reach the theoretical lower bound.
This result (7) is further improved by Cai et al. [26] and Xu et al. [28]. In particular, Cai et al. [26] proposed Truncated Power method (TPM) to remove the requirements on the minimum magnitude of non-zeros (i.e., (6)). Xu et al. [28] introduce an exponential spectrum and proposed Exponential Spectral Pursuit (ESP) algorithm to eliminate the $\log^{3}n$ term in (7). Finally, the sample complexity is reduced to
$m=\Omega\left(ks(1)\log n\right)$ (8)
samples are sufficient for ESP to produce a good estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ without any assumption on $x_{(k)}$. However, this result only achieves the optimal bound (2) for a restrictive class of signals, i.e., those with a single dominant entry where $s(1)=\Theta(1)$. This limitation motivates the development of algorithms that can achieve optimal sample complexity for a broader class of signals.
In this paper, we develop the generalized Exponential Spectral Pursuit (gESP) algorithm to address this challenge. Specifically, gESP selects an index set rather than a single index in its first step (i.e., Step 1 in Table 1). This key modification allows gESP to achieve the optimal sample complexity for a much broader class of signals. Specifically, it relaxes the condition for optimality from the restrictive $s(1)=\Theta(1)$ to the more general $s(\lceil\sqrt{k}\rceil)=\Theta(1)$. We show that when the signal structure $s(p)$ is known, gESP produces a $\delta$-neighborhood estimate with
$m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right)$ (9)
samples. As will be detailed in Section IV, this result is uniformly better than (8).
An alternative approach is also provided if we have no access to $s(p)$; see Corollary 3. In this case, the sample complexity is
$m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$ (10)
It can be shown that in the case where $s(1)=\Theta(\sqrt{k})$ and $s(\sqrt{k})=\Theta(1)$, this result will also be better than (8). The detailed discussion can be found in Section IV.
The remainder of this paper is organized as follows. In Section II, we review the development of the initialization stage and introduce our algorithm. Detailed interpretations are provided step by step. Section III contains theoretical results for the sample complexity of gESP in different cases and makes simple comparisons between these results. In Section IV, discussions are directed in several aspects to illustrate the superiority of our result. In Section V, the proofs of the main theorems are given based on the analysis for each step of gESP. Numerical simulations and analysis are conducted in Section VI. Finally, we conclude our paper in Section VII.
II Algorithms
We introduce some useful notations. Throughout the paper, we use lowercase and uppercase boldface letters to represent vectors and matrices, while employing normal font for real numbers. Let $[n]$ denote the set $\{1,2,3,\cdots n\}$. We denote $x_{(i)}$ as the $i$-largest entry of $\mathbf{x}$ in magnitude. For the sake of readability, we denote $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ as the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. For any set $\subset[n], S{\mathbb{C}}^{S}$ is defined as the subspace of ${\mathbb{C}}^{n}$ spanned by vectors supported on $S$, i.e., $\{\mathbf{x}|\mathbf{x}\in{\mathbb{C}}^{n},{\text{supp}(\mathbf{x})}=S\}$. Unless otherwise specified, we define $\mathbf{a}_{S}$ as the vector that keeps the entries of $\mathbf{a}$ indexed by $S$ while setting others to zero. For the matrix $\mathbf{A}$, we define $\mathbf{A}_{S}$ as the matrix which keeps columns and rows indexed by $S$ while setting others to zero. Given a vector $\mathbf{x}$, the conjugate transpose, $\ell_{2}$ norm and $\ell_{0}$ norm of $\mathbf{x}$ are denoted as $\mathbf{x}^{*}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{0}$ respectively. An $n$-dimensional standard complex Gaussian random vector, denoted as $\mathbf{a}\sim\mathcal{CN}(n)$, is defined as $\mathbf{a}=\mathbf{a}_{1}+\mathrm{j}\mathbf{a}_{2}$, with $\mathbf{a}_{1},\mathbf{a}_{2}\sim\mathcal{N}(\mathbf{0},\frac{1}{2}\mathbf{I})$, and j is the imaginary unit. We use $a\sim b$ to represent that $a$ differs from $b$ by a constant factor. Following the convention in the literature, we use standard asymptotic notations to characterize the complexity as bellow.
  • Big-Omega Notation ($\Omega$): We say $f(n)=\Omega(g(n))$ if there exist positive constants $c$ and $n_{0}$ such that for all $n\geq n_{0}$, we have $0\leq c\cdot g(n)\leq f(n)$. This denotes an asymptotic lower bound.
  • Big-Theta Notation ($\Theta$): We say $f(n)=\Theta(g(n))$ if there exist positive constants $c_{1},c_{2},$ and $n_{0}$ such that for all $n\geq n_{0}$, we have $0\leq c_{1}\cdot g(n)\leq f(n)\leq c_{2}\cdot g(n)$.
Before introducing the proposed algorithm, we first review the prior work. In this section, we temporarily assume that
$\left\lvert\mathbf{x}_{(k)}\right\rvert=\Omega\left(\frac{\|\mathbf{x}\|}{\sqrt{k}}\right).$ (11)
This assumption is made for the convenience of our brief introduction. One can find from our theorems that this assumption is actually unnecessary; also see the results in [26, 28].
II-A Prior work
As introduced in Section I, the initialization stage aims to produce a desired estimate that falls into the $\delta$-neighborhood of the target signal $\mathbf{x}$.
  1. 1.
    Earlier researches contain two steps for initialization.
    Step 1: Construct a spectrum $\mathbf{Z}$ and take the diagonal entries to be the set $\{Z_{j}\}_{j=1}^{n}$. Sorting $\{Z_{j}\}_{j=1}^{n}$ and taking the first $k$ indices yields the final support estimate $S$.
    Generally, the expectation of $\mathbf{Z}$ takes the form as follow
    $\mathbb{E}[\mathbf{Z}]=\alpha\|\mathbf{x}\|^{2}\mathbf{I}+\beta\mathbf{x}\mathbf{x}^{*}.$ (12)
    Thus, the expectation of $Z_{j}$ takes the following form
    $\mathbb{E}[Z_{j}]=\alpha\|\mathbf{x}\|^{2}+\beta|x_{j}|^{2},$ (13)
    where $\alpha,\beta$ are both numerical constants, determined by the specific definition of $Z_{j}$. For instance, [20, 29] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}|\mathbf{a}_{ij}|^{2}$ and the corresponding $\alpha=1,\beta=2$, whereas [14] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}|\mathbf{a}_{ij}|$ and the corresponding $\alpha=\frac{2}{\pi\|\mathbf{x}\|},\beta\approx\frac{1}{6\|\mathbf{x}\|}$. It can be seen that there exists an apparent gap between $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})}$ and $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})^{c}}$. Specifically,
    $\min_{j\in\operatorname*{supp}(\mathbf{x})}\mathbb{E}[Z_{j}]-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}\mathbb{E}[Z_{j}]=$ $\beta|{x_{(k)}}|^{2}$
    $\sim$ $\frac{\|\mathbf{x}\|^{2}}{k}.$ (14)
    Thus, employing some concentration inequalities on $Z_{j}$ and $\mathbb{E}[Z_{j}]$ finally leads to
    $\min_{j\in\operatorname*{supp}(\mathbf{x})}Z_{j}>\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}Z_{j},$ (15)
    which means one can obtain the estimate of support set by sorting the elements of $\{Z_{j}\}_{j=1}^{n}$ and selecting the indices corresponding to the $k$ largest values. Then it can be guaranteed that this support estimate contains all the true candidates. Theoretical analysis shows that deriving an appropriate concentration from the gap $\frac{\|\mathbf{x}\|^{2}}{k}$ (see (14)) requires the sample complexity being [^1] When $Z_{j}$ is statistically heavy-tailed ((e.g., involving the forth power of Gaussian), some extra term containing $\log^{3}n$ (e.g., see (7)) will appear in the complexity. This is because one needs to use a truncation argument to analyze the tail bound. For simplicity, we will ignore such term throughout this subsection.
    $m=\Omega(k^{2}\log n),$ (16)
    which results from the re-scaled gap $\frac{|{x_{(k)}}|^{2}}{\|\mathbf{x}\|^{2}}\leq\frac{1}{k}$.
    Step 2: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    The maximal eigenvector of $\mathbb{E}[\mathbf{Z}]$ is $\mathbf{x}$ and $\mathbb{E}[\|\mathbf{z}\|^{2}]=\|\mathbf{x}\|^{2}$. Therefore, we shall obtain a good estimate of $\mathbf{x}$. Theoretical analysis shows that this step requires
    $m=\Omega(k\log n)$ (17)
    samples. Combining (16) and (17), the final sample complexity is
    $m=\Omega(k^{2}\log n).$ (18)
  2. 2.
    Recently, a novel initialization paradigm is proposed and developed in [27, 26, 28], which contains four steps. Instead of directly estimating the support in one step, an index $j_{\text{max}}$ will be firstly selected, which corresponds to the largest value among $Z_{j}$’s.
    Step 1: Construct a spectrum $\mathbf{Z}$ and take the diagonal entries to be the set $\{Z_{j}\}_{j=1}^{n}$. Select the index $j_{\text{max}}$ corresponding to the largest value among $Z_{j}$’s.
    This step is proved to guarantee $x_{j_{\text{max}}}\sim{x_{(1)}}$ when
    $m=\Omega\left(\frac{\|\mathbf{x}\|^{4}}{|{x_{(1)}}|^{4}}\log n\right).$ (19)
    Step 2: Construct a unit vector $\mathbf{e}_{j_{\text{\rm max}}}$, whose values take $1$ at $j_{\text{max}}$ and 0 at others.
    Step 3: Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{Z}\mathbf{e}_{j_{\text{{\rm max}}}}$ in modulus. [^2] In previous literature, this step and the subsequent one (i.e., Steps 2 & 3) were typically merged into a single step: generating an index set $S^{1}$ corresponding to the largest $k$ entries of the $j_{\text{max}}$ column of $\mathbf{Z}$. Here, we have separated them to align with the structure of our proposed algorithm (Algorithm 1).
    The expectation of $\mathbf{Z}$ still satisfies (12). Thus, we have
    $\mathbf{Z}\mathbf{e}_{j_{\text{max}}}$ $\approx\mathbb{E}[\mathbf{Z}]\mathbf{e}_{{j_{\text{max}}}}$
    $=\alpha\|\mathbf{x}\|^{2}\mathbf{e}_{j_{\text{max}}}+\beta{x_{(1)}}\mathbf{x}.$ (20)
    Establishing some concentration inequalities on this random variable yields the similar result to (15)
    $\min_{j\in\operatorname*{supp}(\mathbf{x})}|\mathbf{Z}\mathbf{e}_{j_{\text{max}}}|_{j}-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}|\mathbf{Z}\mathbf{e}_{j_{\text{max}}}|_{j}\approx\$$\beta|{x_{(1)}}{x_{(k)}}|$
    $\sim\$$\frac{|{x_{(1)}}|\|\mathbf{x}\|}{\sqrt{k}}.$ (21)
    To identify the true support set for (21), it requires
    $m=\Omega\left(k\frac{\|\mathbf{x}\|^{2}}{|{x_{(1)}}|^{2}}\log n\right).$ (22)
    Step 4: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    This step is as the same as Step 2 of the previous method.
    Finally, the overall sample complexity of this method is
    $m=\Omega\left(k\frac{\|\mathbf{x}\|^{2}}{|{x_{(1)}}|^{2}}\log n\right).$ (23)
    Since $\frac{\|\mathbf{x}\|^{2}}{|{x_{(1)}}|^{2}}\leq k$, the enhancement of this method lies in the enlarged gap (see (21)) $\frac{|{x_{(1)}}|\|\mathbf{x}\|}{\sqrt{k}}$, which is potentially larger than $\frac{\|\mathbf{x}\|^{2}}{k}$.
II-B Generalized Exponential Spectral Pursuit
As has been clarified, the enlarged gap $\frac{|{x_{(1)}}|\|\mathbf{x}\|}{\sqrt{k}}$ leads to the improvement in complexity. Intuitively, the larger the gap is, the smaller the sample complexity will be. Note that the term $|{x_{(1)}}|$ in this gap arises from the fact that we simply select one maximal index in Step 1. If more significant indices are chosen in Step 1, the final gap will indeed be enlarged, and thus decrease the complexity in Step 3333In fact, the complexity in Step 1 will increase while that in Step 3 will decrease as more significant indices are chosen. This means that there is a trade-off in the number of the indices that should be chosen in Step 1.. Then the remaining issue is to design Step 2. Assume that we obtain the set $S^{0}$ from Step 1. Following the previous method, a straightforward idea is to construct the vector $\mathbf{e}^{0}$ such that it takes $1$ on $S^{0}$ and 0 for the rest. A more sophisticated design, however, is to compute the maximal eigenvector of $\mathbf{Z}_{S^{0}}$ as $\mathbf{e}^{0}$. This also matches the previous method since $\mathbf{e}_{j_{\text{max}}}$ is exactly the maximal eigenvector of $\mathbf{Z}_{\{j_{\text{max}}\}}$.
So far, we have completed the introduction of the modification idea. The specific algorithm is summarized in Algorithm 1. Now we will further clarify the effectiveness intuitively. The more detailed sample complexity and numerical conclusions are formally presented in Section III.
The spectrum in our proposed gESP algorithm is
$\mathbf{Z}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right)\mathbf{a}_{i}\mathbf{a}_{i}^{*},$ (24)
where $\lambda^{2}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$. It can be shown that $\mathbf{Z}$ satisfies
$\mathbb{E}[\mathbf{Z}]\approx\alpha\|\mathbf{x}\|^{2}\mathbf{I}+\beta\mathbf{x}\mathbf{x}^{*}.$ (25)
and the diagonal entries $Z_{j}$’s satisfies
$\mathbb{E}[Z_{j}]\approx\alpha\|\mathbf{x}\|^{2}+\beta|x_{j}|^{2},$ (26)
Step 1 ensures that
$\|\mathbf{x}_{S^{0}}\|^{2}\sim\sum_{i=1}^{p}|x_{(i)}|^{2}.$ (27)
Since the maximal eigenvector of $\mathbb{E}[\mathbf{Z}_{S^{0}}]$ is $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$, we can approximate $\mathbf{e}_{0}$ as $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$. Then, Step 2 ensures that
$|\mathbf{x}^{*}\mathbf{e}_{0}|=|\mathbf{x}_{S^{0}}^{*}\mathbf{e}_{0}|\approx\|\mathbf{x}_{S^{0}}\|\sim\sqrt{\sum_{i=1}^{p}|x_{(i)}|^{2}}.$ (28)
Note that
$\mathbf{Z}\mathbf{e}^{0}\approx\mathbb{E}[\mathbf{Z}]\mathbf{e}^{0}=\alpha\|\mathbf{x}\|^{2}\mathbf{e}^{0}+\beta\mathbf{x}^{*}\mathbf{e}^{0}\cdot\mathbf{x}.$ (29)
Similar to (21), we have
$\min_{j\in\operatorname*{supp}(\mathbf{x})}|\mathbf{Z}\mathbf{e}^{0}|_{j}-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}|\mathbf{Z}\mathbf{e}^{0}|_{j}$
$\approx$~ $\beta|\mathbf{x}^{*}\mathbf{e}^{0}|\cdot{x_{(k)}}\sim\frac{\sqrt{\sum_{i=1}^{p}|x_{(i)}|^{2}}\|\mathbf{x}\|}{\sqrt{k}}.$ (30)
Therefore, Step 3 along with Step 4 ensure that $\mathbf{z}$ is a good initialization of the target signal $\mathbf{x}$.
Algorithm 1  Generalized Exponential Matching Pursuit
Input: $\mathbf{A}$, $\mathbf{y}$, the sparsity $k$, the parameter pp.
Step 1:
Generate an index set $S^{0}$ corresponding to the largest $p$ diagonal elements of $\mathbf{Z}$.
Step 2: Set $\mathbf{e}^{0}$ as the unit maximal eigenvector of $\mathbf{Z}_{S^{0}}$.
Step 3:
Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{f}=\mathbf{Z}\mathbf{e}^{0}$ in modulus.
Step 4:
Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
Output: $\mathbf{z}$.
III Theoretical Analysis
III-A Preliminary
To begin with, we define the distance between two complex vectors.
Definition 1.
For any $n$-dimensional complex vectors $\mathbf{u},\mathbf{v}\in{\mathbb{C}}^{n}$, the distance between $\mathbf{u}$ and $\mathbf{v}$ is defined as
$\operatorname*{dist}(\mathbf{u},\mathbf{v})=\min_{\phi\in[0,2\pi)}\|\mathbf{u}-e^{\mathrm{j}\phi}\mathbf{v}\|,$ (31)
where $\mathrm{j}$ is the imaginary unit.
In the initialization stage, typical algorithms such as HWF, SPARTA and TP usually aim at producing a good estimate $\mathbf{z}$ falling into the $\delta$ neighborhood of the target signal, i.e.,
$\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|,$ (32)
which is the prerequisite of the refinement stage [19, 20, 29]. In the analysis of gESP, we follow the same line and focus on obtaining the estimate satisfying (32).
Furthermore, we give the definition of the structure function $s(p)$ of the target signal $\mathbf{x}$.
Definition 2.
Let $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ be the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. the structure function $s(p)$ is given by
$s(p)=\frac{\|\mathbf{x}\|^{2}}{\sum_{j=1}^{p}|x_{(j)}|^{2}},~~p=1,2,\cdots,n.$ (33)
The structure function, $s(p)$, quantifies the structural information of a signal $\mathbf{x}$ by measuring the energy distribution across its components. Specifically, $s(p)$ represents the proportion of the total signal energy captured by the $p$ largest magnitude components. Therefore, the behavior of $s(p)$ as $p$ varies directly reflects how energy is concentrated or dispersed within the signal, thus characterizing its structure.
The function $s(p)$ has two important properties that are instrumental for the subsequent analysis.
  1. 1.
    $s(p)$ is monotonously decreasing with respect to $p$. Furthermore,
    $1\leq s(p)\leq\frac{k}{p}.$ (34)
  2. 2.
    Note that
    $ps(p)=\frac{\|\mathbf{x}\|^{2}}{\frac{1}{p}\sum_{i=1}^{p}|x_{(i)}|^{2}},$ (35)
    which implies $ps(p)$ is monotonously increasing with respect to $p$. Furthermore,
    $p\leq ps(p)\leq k.$ (36)
III-B Theoretical results for sample complexity of gESP
In this subsection, we give the theoretical results for Algorithm 1. The corresponding proofs are deferred to Section V.
First, we present the theoretical result for Algorithm 1, and provide a specific value of $p$ when $s(p)$ is known.
Theorem 1.
For any numerical constant $0<\delta<1$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ with probability at least $1-4e^{-c_{\delta}m}$. The corresponding required sample complexity depends on an input parameter $p\in[k]$; for a general $p$, it is given by
$m=\Omega\left(\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right).$ (37)
To achieve the most efficient sampling, this complexity can be minimized by selecting the optimal parameter $p_{\text{opt}}$ that solves
$p_{\text{opt}}={\arg\min}_{p\in[k]}\max\left\{p^{2}s^{2}(p),ks(p)\right\}.$ (38)
By employing this optimal parameter, the minimal sample complexity for the algorithm becomes
$m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right).$ (39)
Remark 1.
The state-of-the-art result [28] is
$m$ $=\Omega\left(\max\left\{s^{2}(1)\log n,ks(1)\log n\right\}\right)$
$=\Omega\left(ks(1)\log n\right).$ (40)
Therefore, our result (39) exhibits a significant improvement.
Remark 2.
To make (39) achieve the theoretical sampling minimum $\Omega(k\log n)$, the following inequalities must hold for some $p\in[k]$:
$s(p)\leq\Theta(1),$ (41)
$p^{2}s^{2}(p)\leq\Theta(k).$ (42)
Since $s(p)\geq 1$ (see (34)), (41) implies $s(p)=\Theta(1)$ for some $p\in[k]$. Then, (42) requires $p\leq c\lceil\sqrt{k}\rceil$ for some constant $c$. Moreover, if $s(p)=\Theta(1)$ holds for some fixed $p\leq c\lceil\sqrt{k}\rceil$, then $s(\lceil\sqrt{k}\rceil)=\Theta(1)$ since $s(p)$ is monotonously decreasing and larger than $1$. Therefore, to achieve the information-theoretical bound, gESP requires $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
This result provides the general sample complexity for gESP. However, in the case where $s(p)$ is unknown, determining the optimal $p$ is challenging. One compromise is to execute Algorithm 1 for $k$ times with $p$ taking each value in $[k]$ in turn. Then we shall obtain a set containing $k$ estimates of the target signal $\mathbf{x}$. It can be guaranteed that at least one estimate falls into the $\delta$-neighborhood of $\mathbf{x}$ with the same sample complexity requirement as (39). The formal statement is in the following corollary.
Corollary 1.
Execute Algorithm 1 for $k$ times with $p$ taking each value in $[k]$ in turn, and collect all the estimates into a set $\mathcal{X}$. Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, there exists at least one estimate $\mathbf{z}\in\mathcal{X}$ that satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ when
$m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p)\log n,ks(p)\log n\right\}\right).$ (43)
When multiple executions of the algorithm with varying $p$ are impractical due to computational constraints, we are compelled to explore the sample complexity for specific, analytically tractable choices of $p$. A natural starting point for such an analysis is the intuitive choice of setting $p=k$. While this can be viewed as a special case of Theorem 1, we demonstrate that a more refined analysis allows us to derive a sharper result, as presented in the following corollary.
Corollary 2.
Take $p=k$ as the input of Algorithm 1. Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number of samples $m$ satisfies
$m=\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right).$ (44)
One may note that this result exactly coincides with the value in (37) with $p=\lceil\sqrt{k}\rceil$. Interestingly, by setting $p=\lceil\sqrt{k}\rceil$ as the input, we can obtain a better result, as demonstrated in the following theorem.
Corollary 3.
Take $p=\lceil\sqrt{k}\rceil$ as the input of Algorithm 1. Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number $m$ of samples satisfies
$m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$ (45)
One one hand, the result (45) cannot be worse than (44) since
$~\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right)$
$\leq$ $~\Omega\left(\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\Big|_{p=\lceil\sqrt{k}\rceil}\log n\right)$
$=$ $~\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right).$ (46)
Moreover, there exist some signals that make the less-than sign strictly holds. Therefore, (45) is indeed tighter than (44). See Section IV-B for an in-depth discussion.
On the other hand, compared to the state of the art, the results (44) and (45) are superior for certain signals. In other words, even though $s(p)$ is unknown as well as executing Algorithm 1 for multiple times is infeasible, our algorithm taking $p=k$ or $\lceil\sqrt{k}\rceil$ can perform better compared with existing algorithms. Discussion IV-A gives an in-depth discussion.
IV Discussion
IV-A Broader Conditions for Optimal sample complexity
As has been discussed in Remark 1 and Remark 2, gESP makes a significant improvement over the state-of-the-art result, if the structural function $s(p)$ is known. This subsection further explores this superiority, demonstrating that gESP achieves the optimal sample complexity for a broader class of signals.
In particular, gESP achieves the information-theoretical bound $\Omega(k\log n)$ when $s(\lceil\sqrt{k}\rceil)=\Theta(1)$, while the state-of-the-art result
$m=\Omega\left(ks(1)\log n\right)$ (47)
requires $s(1)=\Theta(1)$, which is a more strict condition. The improvement results from taking $p$ entries in Step 1 into consideration, thus relaxing the condition on $x_{(1)}$ to $x_{(1)},x_{(1)},\dots,x_{(p)}$. Fig. 1 gives a comprehensive illustration.
Figure 1: An illustration of the relaxed condition for achieving the optimal sample complexity $\Omega(k\log n)$. While state-of-the-art methods require the signal’s energy to be concentrated in its single largest entry: $s(1)=\Theta(1)$, gESP achieves optimality for a broader class of signals where the energy is concentrated among the top $\lceil\sqrt{k}\rceil$ entries: $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
Moreover, even in the case where $s(p)$ is unknown, our results (Corollary 2 and 3) can also perform better in many cases. In the following, we set $p=k$ and consider Corollary 2 for a simple discussion, although we can prove the strict superiority of the result derived in Corollary 3 (see Section IV-B). The result in Corollary 2 is
$m=\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right),$ (48)
Therefore, (48) outperforms the state-of-the-art result (47) once $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$. Here, we offer an example to demonstrate the existence of such signals.
Example 1.
Assume that both $\sqrt{k}$ and $\sqrt[6]{k}$ are integers [^4] In fact, we can also construction such examples even though this assumption does not hold. However, omitting it will lead to unnecessary complications in the construction (e.g., involving terms like $\lceil\sqrt{k}\rceil\pm 1$). The same phenomenon also occurs in Example 2. For the sake of simplicity, we make this assumption. The target signal $\mathbf{x}$ satisfies
$|x_{(i)}|^{2}=\begin{cases}\frac{\|\mathbf{x}\|^{2}}{\sqrt{k}},&i=1,\\ \frac{\|\mathbf{x}\|^{2}}{\sqrt{k}-1}\times\left(\frac{1}{\sqrt[6]{k}}-\frac{1}{\sqrt{k}}\right),&i=2,\cdots,\sqrt{k},\\ \frac{\|\mathbf{x}\|^{2}}{k-\sqrt{k}}\times\left(1-\frac{1}{\sqrt[6]{k}}\right),&i=\sqrt{k}+1,\cdots,k.\end{cases}$
In this case, $s(1)=\sqrt{k}$ and $s(\sqrt{k})=\sqrt[6]{k}$, and one can thus find (48) outperforms (47).
Interestingly, when the condition $s(\lceil{\sqrt{k}}\rceil)=\Theta(1)$ holds, our result (48) also achieves the information-theoretic bound (2). Significantly, even without prior knowledge of the signal structure $s(p)$, this optimality is attainable for the same class of signals in the case $s(p)$ is known, simply by setting the parameter $p=k$.
IV-B Strict superiority of Corollary 3 over Corollary 2
In Corollary 3, we claim that the result
$m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right),$
is better than (48) in Section III-B. In this section, we provide one certain example to illustrate this strict superiority.
Example 2.
Assume that both $\sqrt{k}$ and $\sqrt[4]{k}$ are integers. Construct the target signal $\mathbf{x}$ as
$|x_{(i)}|^{2}=\begin{cases}\frac{\|\mathbf{x}\|^{2}}{\sqrt[4]{k^{3}}},&i=1,2,\cdots,\sqrt[4]{k},\\ \frac{\|\mathbf{x}\|^{2}}{\sqrt{k}-\sqrt[4]{k}}\times\left(\frac{1}{\sqrt[3]{k}}-\frac{1}{\sqrt{k}}\right),&i=\sqrt[4]{k}+1,\cdots,\sqrt{k},\\ \frac{\|\mathbf{x}\|^{2}}{k-\sqrt{k}}\times\left(1-\frac{1}{\sqrt[3]{k}}\right),&i=\sqrt{k}+1,\cdots,k.\end{cases}$
In this case, it can be calculated that the result in Corollary 2 is
$\Omega\left(ks^{2}(\sqrt{k})\log n\right)=\Omega(k^{5/3}\log n),$
while that in Corollary 3 is
$\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right)$
$=$ $~\Omega\left(ks(\sqrt[4]{k})\log n\right)=\Omega(k^{3/2}\log n).$
This demonstrate the improvement in Corollary 3.
V Proof
The proofs of Theorem 1 and Corollary 13 are mainly based on the analysis for each step of gESP. In the following, we provide several propositions with respect to each step respectively.
  • Step 1: Generate an index set $S^{0}$ corresponding to the largest $p$ diagonal elements of $\mathbf{Z}$.
    Proposition 1.
    When the number of samples satisfies $m\geq Cp^{2}s^{2}(p)\log n$, it holds with probability exceeding $1-2e^{-cm}$ that
    $\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)},$ (49)
    where $c$ and $C$ are universal constants.
  • Step 2: Set $\mathbf{e}^{0}$ as the unit maximal eigenvector of $\mathbf{Z}_{S^{0}}$.
    Proposition 2.
    Suppose the set $S^{0}$ contains $p^{\prime}$ indices and satisfies (49), then for any constant $0<\mu_{1}<1$, the inequality
    $|\mathbf{x}^{*}\mathbf{e}^{0}|\geq\mu_{1}\|\mathbf{x}_{S^{0}}\|$ (50)
    holds with probability at least $1-2e^{-c_{\mu_{1}}m}$ with the proviso that $m\geq{C_{\mu_{1}}}{p^{\prime}s^{2}(p)}\log n$, where $c_{\mu_{1}}$ and $C_{\mu_{1}}$ are constants related to $\mu_{1}$.
    Remark 3.
    It may seem unusual to introduce a new parameter $p^{\prime}$ in this proposition, given that $S^{0}$ contains $p$ indices in Algorithm 1. However, the proposition holds even when $p^{\prime}=p$, making it a generalized version. We emphasize that this generalization will be useful in our proofs of Corollary 2 and 3, where we do not have access to the optimal choice of $p$.
  • Step 3: Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{f}=\mathbf{Z}\mathbf{e}^{0}$ in modulus.
    Proposition 3.
    Denote the set $S_{\gamma}$ as
[…]
    where $\gamma\in(0,1)$ is a constant. Moreover, suppose that (49) holds and $\mathbf{e}^{0}$ satisfies (50). Then, if the number of samples satisfies $m\geq C_{\mu_{1}}\gamma^{-2}{ks(p)}\log n$, it holds with probability at least $1-4e^{-c_{\mu_{1},\gamma}m}$ that
    $S^{1}\supseteq S_{\gamma},$
    where $\mu_{1}$ is the hyperparameter given in (50) and $c_{\mu_{1},\gamma}$ is the constant depending on $\mu_{1}$ and $\gamma$. In other words, gESP shall select all the indices of $S_{\gamma}$ in Step 3.
  • Step 4: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
    Proposition 4.
    Suppose $S^{1}\supseteq S_{\gamma}$. For any constant $0<\delta<1$ satisfying
    $\delta^{2}\geq 2-2\sqrt{\left(1-\frac{\gamma^{2}}{4}\right)^{3}}+\frac{\gamma^{2}}{4},$ (51)
    Step 4 produces a signal estimation $\mathbf{z}$ falling into the $\delta$-neighborhood of $\mathbf{x}$, i.e.,
    $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ (52)
    with probability exceeding $1-2e^{-c_{\delta}m}$ providing that $m\geq{C_{\delta}}{k}\log n$, where $c_{\delta}$ and $C_{\delta}$ are numerical constants depending on $\delta$.
Now we are proceeding to prove the theorems in Section III.
Figure 2: An Illustration of the two functions in (56).
Proof of Theorem 1.
Collecting all the results in Proposition 14 with $p^{\prime}=p$, we can show that the sample complexity is
$m=\Omega\left(\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right),$ (53)
and the overall probability is $1-e^{-cm}$. Since we take the optimal value for $p$ as
$p_{\text{opt}}={\arg\min}_{p\in[k]}\max\left\{p^{2}s^{2}(p),ks(p)\right\},$
we obtain the desired results. ∎
Proof of Corollary 1.
For each fixed $p\in[k]$, collecting all the results in Proposition 14 with $p^{\prime}=p$, we can show that the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-e^{-cm}$ when
$m=\Omega\left(\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right).$
To ensure that at least one estimate $\mathbf{z}$ satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$, we can apply the union bound for all $p\in[k]$. Then the sample complexity can be reduced to
$m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p)\log n,ks(p)\log n\right\}\right),$
and the probability is at least $1-ke^{-cm}$. It is actually $1-e^{-cm}$ if we take a sufficiently large constant $c$ since $m>k$. Thus the proof is complete. ∎
(a) Gaussian signal. $k$ = 10.
(b) 0-11 signal. $k$ = 10.
(c) Exponential decaying signal. $k$ = 10
Figure 3: Performance comparison of relative error and fraction of recovered support as a function of sampling ratio.
Proof of Corollary 2.
For any $p\in[k]$, denote $S^{\prime}$ as the set corresponding to the largest $p$ diagonal elements of $\mathbf{Z}$. Note that $S^{0}$ corresponds to the largest $k$ diagonal elements of $\mathbf{Z}$. It holds that
$\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{\|\mathbf{x}\|^{2}}\overset{(a)}{\geq}\frac{\|\mathbf{x}_{S^{\prime}}\|^{2}}{\|\mathbf{x}\|^{2}}\overset{(b)}{\geq}\frac{1}{2s(p)},$ (54)
where $(a)$ is because $k\geq p$, and $(b)$ holds via applying Proposition 1 on the set $S^{\prime}$ when $\geq Cp^{2}s^{2}(p)\log n$.
Then, from the results above and in Proposition 24 (with $p^{\prime}=k$), it holds that for any $p\in[k]$, when
$m=\Omega\left(\max\left\{p^{2}s^{2}(p)\log n,ks^{2}(p)\log n\right\}\right),$ (55)
the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-e^{-cm}$. This result also holds for all $p\in[k]$ since we can take the union bound, and the probability will be $1-ke^{-cm}$. It is actually $1-e^{-cm}$ if we take a sufficiently large constant $c$ since $m>k$.
Note that the choice of $p\in[k]$ is arbitrary. Therefore, we can choose the best $p$ to make the sample complexity (55) smallest, which becomes
$m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p)\log n,ks^{2}(p)\log n\right\}\right),$ (56)
The remaining thing is to show that (56) is equivalent to (44).
As stated following Definition 2, $ps(p)$ is monotonously increasing, and $s(p)$ is monotonously decreasing. Both of them are positive. Accordingly, we can make an illustrative plot for the two functions in (56); see Fig. 2.
Therefore, the minimum occurs when
$p=\lceil{\sqrt{k}}\rceil~~~~\text{or}~~~~p=\lceil{\sqrt{k}}\rceil-1.$ (57)
From Lemma 1, we know that both the cases imply the same sample complexity
$m=\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right),$ (58)
which completes the proof.
Proof of Corollary 3.
The proof is similar to the proof of Corollary 2. Assume any fixed $p\in[\lceil{\sqrt{k}}\rceil]$. The sample complexity for Step 1 is still $\Omega(p^{2}s^{2}(p)\log n)$, and applying Proposition 24 on the rest steps with $p^{\prime}=\lceil{\sqrt{k}}\rceil$ finally leads to
$m=\Omega\left(\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$ (59)
Since the choice of $p\in[k]$ is arbitrary, taking the union bound yields the desired result
$m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$
VI Numerical Simulations
In this section, we present numerical experiments to evaluate the performance of our proposed gESP algorithm against several state-of-the-art methods.
VI-A Experimental Setup
Across all simulations, the signal dimension is fixed at $n=1000$. The measurement vectors $\{\mathbf{a}_{i}\}_{i=1}^{m}\subset\mathbb{C}^{n}$ are i.i.d. standard complex Gaussian vectors. The $k$-sparse signal $\mathbf{x}\in\mathbb{C}^{n}$ is constructed by first selecting a support of size $k$ uniformly at random. The non-zero entries are then generated from one of three models: i) i.i.d. standard complex Gaussian, ii) all ones (binary), and iii) exponentially decaying magnitudes. All reported results are averaged over $1000$ independent trials to mitigate random fluctuations.
We compare gESP with CoPRAM, TPM, and ESP. For a fair comparison, we consider only the initialization stage of CoPRAM, and use the recommended hyperparameters for TPM. For gESP, we set the theoretically optimal parameter $p_{opt}$ from (38) for each trail.
The performance metric is the relative error, defined as:
$\text{Relative Error}=\frac{\|\mathbf{z}-\mathbf{x}\|_{2}}{\|\mathbf{x}\|_{2}},$
where $\mathbf{z}$ is the estimated signal and $\mathbf{x}$ is the true signal [26, 30].
VI-B Performance Comparison
In this experiment, we vary the sampling ratio $m/n$ from $0.05$ to $1.0$ and record the corresponding relative error.
As depicted in Figs. 3a3c, the performance on sparse Gaussian signals provides a direct and strong validation of our theoretical analysis. For this signal class, gESP was configured using the parameter $p$ determined by our theoretically derived optimum. Consistent with the predictions of our theory, gESP achieves the best performance, consistently and slightly outperforming ESP and other state-of-the-art methods. This clear correspondence between our theory and the empirical results underscores the accuracy of our analysis for signals with rich amplitude variations.
In contrast, the results for binary and exponentially decaying signals, while not perfectly aligned with asymptotic predictions, offer valuable insights into the practical behavior of the algorithms. For binary signals, theory suggests the complexity would always be $\Omega(k^{2}\log n)$ no matter what value $p$ is, and thus one might expect similar performance across methods. However, we set $p=k$ in gESP, and it demonstrates a surprisingly large performance margin. We conjecture that this significant practical advantage stems from the influence of non-asymptotic constant factors that our analysis does not model but which clearly favor the gESP framework. Similarly, for exponentially decaying signals, gESP is marginally outperformed by ESP. This is also likely an effect of such constants; our theory provides the optimal scaling for $p$, but in this finite-dimensional regime, the true optimal value is empirically found to be $p=1$ (i.e., ESP). This does not contradict our theory but rather highlights that while it provides a powerful general guideline, the precise optimal parameter in practice can be influenced by the signal’s specific structure.
VII Conclusion
In this paper, we investigate the problem of sparse phase retrieval, focusing on improving initialization algorithms for recovering a $k$-sparse $n$-dimensional signal from $m$ phaseless observations. Through an analysis of existing methods, we propose generalized Exponential Spectral Pursuit (gESP), which enhances initialization performance and refines the understanding of sample complexity in sparse phase retrieval. Empirical evaluations further validate the practical effectiveness of gESP. Simulation experiments demonstrated its robustness and efficiency in recovering sparse signals, highlighting its potential utility in real-world applications.
While this work advances both theoretical bounds and practical performance for sparse phase retrieval, it does not fully resolve the statistical-to-computational gap. Specifically, under Gaussian measurements, $\Omega(k\log n)$ samples are theoretically sufficient to determine the target signal. However, in practical scenarios requiring polynomial-time algorithms, the sample complexity increases to $\Omega(k^{2}\log n)$ in the worst case, such as when the target signal consists of 0-11 entries. Future research could aim to bridge this gap by exploring novel algorithmic frameworks or leveraging insights from other areas, such as computational complexity and optimization. Additionally, extending the applicability of gESP to broader measurement models and refining the refinement stage offer promising directions for advancing this field.
Appendix A Basic tools and Lemmas
In this section, we present several useful lemmas and necessary proofs. First of all, recall that
$\mathbf{Z}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right)\mathbf{a}_{i}\mathbf{a}_{i}^{*},$ (A.1)
where $\lambda^{2}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$. It is challenging to derive the expectation of $\mathbf{Z}$. Following the technique in [12, 28], we introduce
$\tilde{\mathbf{Z}}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{a}_{i}\mathbf{a}_{i}^{*},$ (A.2)
whose expectation was derived in [12] as
$\mathbb{E}[\tilde{\mathbf{Z}}]=\frac{\mathbf{x}\mathbf{x}^{*}}{4\|\mathbf{x}\|^{2}}.$ (A.3)
For convenience, denote
$Z_{j}$ $=\operatorname*{diag}(\mathbf{Z})_{j},$ $\tilde{Z}_{j}$ $=\operatorname*{diag}(\tilde{\mathbf{Z}})_{j},$ (A.4)
$\mathbf{f}$ $=\mathbf{Z}\mathbf{e}^{0},$ $\tilde{\mathbf{f}}$ $=\tilde{\mathbf{Z}}\mathbf{e}^{0}.$
In the following, we introduce several results concentrating on the effect of the ceiling sign on the sample complexity.
Lemma 1.
For the ceiling function in the asymptotic notation, we have
$\Omega(\lceil\sqrt{k}\rceil)$ $=\Omega(\sqrt{k}),$ (A.5)
$\Omega(s(\lceil\sqrt{k}\rceil)$ $=\Omega(s(c\lceil\sqrt{k}\rceil)),$ (A.6)
where $c$ is a fixed positive integer.
Proof.
Note that $\sqrt{k}\leq\lceil\sqrt{k}\rceil\leq\sqrt{k}+1$, then (A.5) can be easily derived from the following relationship.
$\lim_{k\to\infty}\frac{\sqrt{k}+1}{\sqrt{k}}=1.$
As for (A.6), recall from the definition of $s(p)$, we have
$s(c\lceil\sqrt{k}\rceil)=\frac{\|\mathbf{x}\|^{2}}{\sum_{j=1}^{c\lceil\sqrt{k}\rceil}|x_{(j)}|^{2}}.$
Since $x_{(j)}$’s are the rearrangement of the entries of $\mathbf{x}$ with a descend order in their modulus, the summation in the denominator can be estimated as
$\sum_{j=1}^{\lceil\sqrt{k}\rceil}|x_{(j)}|^{2}\leq\sum_{j=1}^{c\lceil\sqrt{k}\rceil}|x_{(j)}|^{2}\leq c\sum_{j=1}^{\lceil\sqrt{k}\rceil}|x_{(j)}|^{2},$
which in turn yields
$\frac{1}{c}s(\lceil\sqrt{k}\rceil)\leq s(c\lceil\sqrt{k}\rceil)\leq s(\lceil\sqrt{k}\rceil).$
In that $c$ is a constant as $k$ and $n$ tend to infinity, (A.6) is a natural derivation. ∎
Lemma 2 (Bernstein inequality for complex cases).
Let $X_{i}$ be independent, mean-zero complex sub-exponential random variables for $i\in[m]$, then for any $\epsilon>0$, we have
$\mathbb{P}\left(\left\lvert\frac{1}{m}\sum_{i=1}^{m}X_{i}\right\rvert\geq\epsilon\right)\leq 4\exp\left(-cm\min\left\{\frac{\epsilon^{2}}{K^{2}},\frac{\epsilon}{K}\right\}\right),$
where $K=\mathop{\max}_{i\in[m]}\|X_{i}\|_{\psi_{1}}$ and $c$ is the numerical constant. Also see [22, Lemma 10].
Next, the difference between $\lambda^{2}$ and $\|\mathbf{x}\|^{2}$ is analyzed.
Lemma 3.
Given $k$-sparse signal $\mathbf{x}$, for any constant $\beta>0$, suppose that $\mathbf{a}_{i}$’s are complex Gaussian random variables. Then provided that $m\geq C_{\beta}\kappa^{2}$,
$\left(1-\frac{\beta}{\kappa}\right)\|\mathbf{x}\|^{2}\leq\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}\leq\left(1+\frac{\beta}{\kappa}\right)\|\mathbf{x}\|^{2}$
holds with probability at least $1-e^{-c_{\beta}m}$, where $\kappa$ can be substitute by any quantity satisfying $\frac{\beta}{\kappa}<1$, such as $ps(p)$, and $c_{\beta}$ and $C_{\beta}$ are numerical constants determined by $\beta$.
Proof.
It is easy to find that $|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}$’s are independent sub-exponential random variables with mean $\mathbb{E}[|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}]=\|\mathbf{x}\|^{2}$ and $c_{0}\|\mathbf{x}\|^{2}$ as their squared sub-exponential norms. Hence, employing Bernstein’s inequality yields
$\mathbb{P}\left(\left\lvert\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}-\|\mathbf{x}\|^{2}\right\rvert>\frac{\beta}{\kappa}\|\mathbf{x}\|^{2}\right)$
$~~~~~~~\leq 2\exp\left(-C_{\beta}m\min\left\{\frac{\beta^{2}}{\kappa^{2}K^{2}},\frac{\beta}{\kappa K}\right\}\right),$
where $K=\mathop{\max}_{i\in[m]}\||\mathbf{a}_{i}^{*}\mathbf{x}|^{2}\|_{\psi_{1}}/\|\mathbf{x}\|^{2}$ and $C_{\beta}$ is the numerical constant. Thus we obtain the desired result.∎
Finally, we introduce a useful lemma that measures the similarity between the exponential spectral $\mathbf{Z}$ and its estimator $\mathbb{E}[\tilde{\mathbf{Z}}$].
Lemma 4.
Suppose a constant $\eta>0$ and a positive integer $q\leq k$. For all $S\subseteq[n]$ satisfying $|S|=q$, let $T$ be a subset of $\mathbb{C}^{n}$ such that $\|\mathbf{z}\|\leq 1$ holds for all $\mathbf{z}\in T$. Given any constant $\epsilon\in(0,1)$, let $\mathcal{N}(T,\epsilon)$ be the covering number [31, Definition 4.2.2] of $T$ and $g(k,s(p))$ be a function with respect to $k$ and $s(p)$ satisfying $g(k,s(p))\geq 1$. It holds with probability at least $1-2e^{-c_{\eta}m}$ that
$\sup_{\mathbf{z}_{1},\mathbf{z}_{2}\in T}\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{2}\leq\frac{\eta}{g(k,s(p))}$ (A.7)
with the condition that
$m\geq C_{\eta}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)),$ (A.8)
where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
In particular, when $T=\{\mathbf{u}|\operatorname{supp}(\mathbf{u})=S,\|\mathbf{u}\|\leq 1\}$, the inequality
$\left\lVert\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right\rVert\leq\frac{\eta}{2g(k,s(p))}$ (A.9)
holds true given $m\geq C_{\eta}qg^{2}(k,s(p))\log n$.
Proof.
Denote the smallest $\epsilon$-net of $T$ as $N_{\epsilon}^{T}$, which means that $|N_{\epsilon}^{T}|=\mathcal{N}(T,\epsilon)$. Then, for any unit $\mathbf{z}_{1},\mathbf{z}_{2}\in T$, there exist $\mathbf{z}_{10}\in N_{\epsilon}^{T}$ and $\mathbf{z}_{20}\in N_{\epsilon}^{T}$ such that $\|\mathbf{z}_{1}-\mathbf{z}_{10}\|\leq\epsilon$ and $\|\mathbf{z}_{2}-\mathbf{z}_{20}\|\leq\epsilon$. It holds that
$\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{2}$
== $\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}+\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)(\mathbf{z}_{2}-\mathbf{z}_{20})$
$+(\mathbf{z}_{1}-\mathbf{z}_{10})^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}$
$\overset{(a)}{\leq}$ $\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}+2\epsilon\left\lVert\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right\rVert,$
where $(a)$ comes from the definition of $N_{\epsilon}^{T}$ and the matrix norm, together with the fact that $\|\mathbf{z}_{20}\|\leq 1$. We first prove the following two inequalities
$\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))},$ (A.10)
$\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))}.$ (A.11)
$\dagger$} Proof for (A.10)
Simple calculation gives that
$\mathbf{z}_{10}^{*}\tilde{\mathbf{Z}}_{S}\mathbf{z}_{20}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{\mathbf{y}_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20},$
which is the sum of independent sub-exponential random variables with constant $\psi_{1}$ norms, since $\|\mathbf{z}_{10}\|,\|\mathbf{z}_{20}\|\leq 1$. Employing Bernstein’s inequality yields
$\mathbb{P}\left(\left\lvert\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\right\rvert>\frac{\eta}{4}\times\frac{1}{g(k,s(p))}\right)$
$\leq$~ $2\exp\left(-C^{\prime}_{\eta}m\min\left\{\frac{\eta^{2}}{16K^{2}g^{2}(k,s(p))},\frac{\eta}{4Kg(k,s(p))}\right\}\right)$
$\doteq$~ $p_{1},$
where the sub-exponential norm satisfies
$K=\max_{i\in[m]}\left\rVert\left(\frac{1}{2}-\exp\left(-\frac{\mathbf{y}_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\left\lvert(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{10}\right\rvert\left\lvert(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert\right\rVert_{\psi_{1}}$
and $c_{\eta}$ is the numerical constant.
Take the union bound for all $\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}$, we have
$~\mathbb{P}\left(\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))}\right)$
$\geq$ $~1-\mathcal{N}^{2}(T,\epsilon)\times p_{1}.$
Furthermore, take the union bound for all $S\subseteq[n]$ satisfying $|S|\leq q$, the probability will exceed $1-{n\choose q}\mathcal{N}^{2}(T,\epsilon)p_{1}$.
Therefore, when
$m\geq C_{\eta}^{\prime}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)),$
(A.10) holds with probability exceeding $1-2\exp\big(c^{\prime}_{\eta}m\big)$.
$\dagger$} Proof for (A.11)
Direct calculation yields
$\left\rvert\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\right)\mathbf{z}_{20}\right\rvert$
== $\left\lvert\frac{1}{m}\sum_{i=1}^{m}\left(\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{z}_{10}^{*}(\mathbf{a}_{i}\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$ (A.14)
$\overset{(a)}{=}$ $\max_{i\in[m]}\exp\left(-\frac{y_{i}^{2}}{\xi}\right)\hskip-0.7113pt\frac{y_{i}^{2}}{\xi^{2}}\hskip-0.7113pt\big|\lambda^{2}-\|\mathbf{x}\|^{2}\big|\hskip-0.7113pt\hskip-0.7113pt\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i}\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$ (A.17)
$\overset{(b)}{\leq}$ $~\frac{1}{\xi}\big|\lambda^{2}-\|\mathbf{x}\|^{2}\big|\times\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$, (A.20)
where $(a)$ comes from Lagrange’s mean value formula with $F(t)=\exp\big(\frac{y_{i}^{2}}{t}\big)$ and $\xi\in\left[\min\{\lambda^{2},\|\mathbf{x}\|^{2}\},\max\{\lambda^{2},\|\mathbf{x}\|^{2}\}\right]$, while $(b)$ employs the numerical inequality $xe^{-x}\leq 1,\ x\geq 0$.
For the second term, since $\mathbf{a}_{i}$’s are complex standard Gaussian random vectors and $\|\mathbf{z}_{10}\|,\|\mathbf{z}_{20}\|\leq 1$, it can be derived that for every $i\in[m]$, $\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}$ is sub-exponential random variable with constant $\psi_{1}$ norm and expectation being $(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}$. Therefore, follow the same concentration analysis as in the proof of (A.10), we conclude that the inequality
$\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}-(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}\right\rvert\leq\frac{\eta}{12}\times\frac{1}{g(k,s(p))}$ (A.21)
holds for all $\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}$ and $S\subseteq[n]$ satisfying $|S|\leq q$ with probability exceeding $1-2\exp(c^{\prime\prime}_{\eta}m)$ when
$m\geq C^{\prime\prime}_{\eta}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)).$
Note that
$\left\lvert(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}\right\rvert\leq\sqrt{\|\mathbf{z}_{10}\|^{2}\|\mathbf{z}_{20}\|^{2}}\leq 1.$
The inequality (A.21) can be transformed to
$\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert\leq 1+\frac{\eta}{12}\times\frac{1}{g(k,s(p))}.$ (A.22)
For the first term, by employing Lemma 3 with with $\kappa=g(k,s(p))$ and $\beta=\eta/12$, we conclude that
$\frac{1}{\xi}\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert$ $\leq\frac{1}{\|\mathbf{x}\|^{2}-\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert}\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert$
$\leq\frac{1}{\left(1-\frac{\eta}{12g(k,s(p))}\right)\|\mathbf{x}\|^{2}}\times\frac{\eta\|\mathbf{x}\|^{2}}{12g(k,s(p))}$ (A.23)
holds true with probability exceeding $1-e^{-cm}$ as soon as
$m\geq C^{\prime\prime}_{\eta}g^{2}(k,s(p)).$
Finally, taking (A.22) and (A.23) into (A.20) yields
[…]
$\overset{(a)}{\leq}$ $~3\times\frac{\eta}{12}\times\frac{1}{g(k,s(p))}=\frac{\eta}{4}\times\frac{1}{g(k,s(p))},$
where $(a)$ comes from the property $\frac{1+x}{1-x}\leq 3$ with $x\leq\frac{1}{2}$ and the fact that $\frac{\eta}{12}\times\frac{1}{g(k,s(p))}\leq\frac{1}{2}$. Here, we complete the proof of two sub-conclusions. Finally, combining (A.10) and (A.11) with $T=\{\mathbf{u}|\operatorname{supp}(\mathbf{u})=S,\|\mathbf{u}\|\leq 1\}$ yields (A.9), which in turn proves our conclusion. ∎
Remark 4.
As indicated above, the term $q\log(n/q)$ arises from the union bound for $S\in[n],|S|=q$. For fixed set $S$, the sample complexity reduces to
$m\geq C_{\eta}g^{2}(k,s(p))\log\mathcal{N}(T,\epsilon).$ (A.24)
A-A Concentration analysis for the sum of the diagonal entries
Lemma 5.
For all $\mathcal{S}\subseteq[n]$ with $|\mathcal{S}|=p$ and constant $0<\eta<1$, the following concentration inequality
$\left\lvert\frac{1}{p}\sum_{j\in\mathcal{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert<\eta\times\frac{1}{ps(p)}$ (A.25)
holds true with probability exceeding $1-2e^{-c_{\eta}m}$ given $m\geq C_{\eta}p^{2}s^{2}(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
Proof.
We can see from (A.3) that
$\mathbb{E}[\tilde{Z}_{j}]=\left\{\begin{array}[]{cc}0,&j\notin{\text{supp}(\mathbf{x})},\\ \frac{|x_{j}|^{2}}{4\|\mathbf{x}\|^{2}},&j\in{\text{supp}(\mathbf{x})}.\end{array}\right$. (A.26)
Therefore, the left term of (A.25) can be expressed as
$\frac{1}{p}\sum_{j\in\mathcal{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)=\frac{1}{p}\sum_{j\in\mathcal{S}}\left(\mathbf{Z}_{\{j\}}\hskip-0.7113pt-\hskip-0.7113pt\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right).$
For all $j\in[n]$, employing Lemma 4 with $S=\{j\}$, the fixed $T=\{\mathbf{1},-\mathbf{1}\}$, and $g(k,s(p))=ps(p)$, we can derive that
$\sup_{j\in[n]}\mathbf{1}^{*}\left(\mathbf{Z}_{\{j\}}-\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right)\mathbf{1}\leq\eta\times\frac{1}{ps(p)},$
$\sup_{j\in[n]}\mathbf{1}^{*}\left(\mathbf{Z}_{\{j\}}-\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right)(-\mathbf{1})\leq\eta\times\frac{1}{ps(p)}.$
holds with probability exceeding $1-2e^{-c_{\eta}m}$ and the corresponding sample complexity becomes
$m\geq C_{\eta}p^{2}s^{2}(p)\log n,$
which in turn proves (A.25) since $\mathcal{S}\subseteq[n]$.
A-B Concentration analysis for $\mathbf{f}$ in (A.4)
Lemma 6.
For any constant $\eta>0$ and $\gamma>0$, the following concentration inequality
$\max_{l\in[n]}\left\lvert f_{l}-\mathbb{E}(\tilde{f}_{l})\right\rvert<\eta\times\frac{\gamma}{2\sqrt{ks(p)}}$ (A.27)
holds with probability exceeding $1-2\exp(-c_{\eta}m)$ given $m\geq C_{\eta}\frac{ks(p)}{\gamma^{2}}\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constant related to $\eta$.
Proof.
Note that the expression of $f_{l}$ can be transformed as
$f_{l}=\mathbf{e}_{l}^{*}\mathbf{Z}_{S^{0}}\mathbf{e}^{0},$
where $\mathbf{e}_{l}\in{\mathbb{C}}^{n}$ is the vector that sets its $l$-th entry as 1 while others as 0, $S^{0}$ is a fixed set derived by Step 1 of gESP and $\mathbf{e}^{0}$ is the principle eigenvector of $\mathbf{Z}_{S^{0}}$. Employing Lemma 4 with fixed $S=S^{0}$, fixed $T=\{\mathbf{e}_{1},\mathbf{e}_{2},\cdots,\mathbf{e}_{n},\mathbf{e}^{0}\}$ and $g(k,s(p))=\frac{2}{\gamma}\sqrt{ks(p)}$, we derive that
$\max_{l\in[n]}\left\lvert f_{l}-\mathbb{E}(\tilde{f}_{l})\right\rvert<\eta\times\frac{\gamma}{2\sqrt{ks(p)}}$
holds true with probability exceeding $1-2\exp(-c_{\eta}m)$ with the sample complexity satisfying $m\geq C_{\eta}\frac{ks(p)}{\gamma^{2}}\log n$.
Appendix B Proof for Proposition 1
Denote $\overline{S}$ as the index set corresponding to the largest $p$ elements of $\mathbf{x}$. We establish the upper and lower bounds for $\frac{1}{p}\sum_{j\in S^{0}}Z_{j}$ and $\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}$, respectively.
$\frac{1}{p}\sum_{j\in S^{0}}Z_{j}\leq$ $~\frac{1}{p}\sum_{j\in S^{0}}\mathbb{E}\left[\tilde{Z}_{j}\right]+\left\lvert\frac{1}{p}\sum_{j\in S^{0}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$\leq$ $~\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{4p\|\mathbf{x}\|^{2}}+\frac{1}{16ps(p)},$ (B.1)
$\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}\geq$ $~\frac{1}{p}\sum_{j\in\overline{S}}\mathbb{E}\left[\tilde{Z}_{j}\right]-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$=$ $\frac{\|\mathbf{x}_{\overline{S}}\|^{2}}{4p\|\mathbf{x}\|^{2}}-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
$\geq$ $~\frac{1}{4ps(p)}-\frac{1}{16ps(p)}=\frac{3}{16ps(p)}.$ (B.2)
From the definition of $S^{0}$ and $\overline{S}$, we have
$\frac{1}{p}\sum_{j\in S^{0}}Z_{j}\geq\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}.$ (B.3)
Therefore, combining (B.1), (B.2) and (B.3), we have
$\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{4p\|\mathbf{x}\|^{2}}+\frac{1}{16ps(p)}\geq\frac{3}{16ps(p)},$ (B.4)
which implies
$\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}.$ (B.5)
Thus we complete the proof.
Appendix C Proof for Proposition 2
For any set $\mathcal{T}$} satisfying $|{\mathcal{T}}|=p^{\prime}\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$, suppose $\mathbf{z}$ as the unit eigenvector corresponding to the largest eigenvalue, denoted as $\tau$, of
$\mathbf{Z}_{{\mathcal{T}}}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right)(\mathbf{a}_{i})_{{\mathcal{T}}}(\mathbf{a}_{i})^{*}_{{\mathcal{T}}}.$
Then, we have
[…]
$\overset{(a)}{\leq}$ $\left\lVert\mathbf{Z}_{{\mathcal{T}}}-\frac{\mathbf{x}_{{\mathcal{T}}}\mathbf{x}^{*}_{{\mathcal{T}}}}{4\|\mathbf{x}\|^{2}}\right\rVert\|\mathbf{z}\|^{2}\overset{(b)}{\leq}\frac{\eta}{s(p)},$ (C.6)
where $(a)$ is based on the relationship between matrix norm and vector norm, $(b)$ employs (A.9) with $g(k,s(p))=s(p)$ and the condition $\|\mathbf{z}\|=1$. The above inequality holds with probability at least $1-2e^{-c_{\eta}m}$ when $m\geq C_{\eta}p^{\prime}s^{2}(p)\log n$. Therefore, through simple transformation, (C.6) can be expressed as
$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\geq 4(\tau-\frac{\eta}{s(p)})\|\mathbf{x}\|^{2}.$ (C.7)
Moreover, since $\tau$ is the largest eigenvalue of $\mathbf{Z}_{{\mathcal{T}}}$, we can estimate it as
[…]
== $~\frac{1}{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}\mathbf{x}^{*}_{{\mathcal{T}}}\left(\mathbf{Z}_{{\mathcal{T}}}-\frac{\mathbf{x}_{{\mathcal{T}}}\mathbf{x}^{*}_{{\mathcal{T}}}}{4\|\mathbf{x}\|^{2}}\right)\mathbf{x}_{{\mathcal{T}}}+\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}}$ (C.13)
$\overset{(a)}{\geq}$ $~-\frac{\eta}{s(p)}+\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}},$ (C.16)
where $(a)$ comes from the definition of matrix norm and Lemma 4 ((A.9)) with g(k,s(p))=s(p)g(k,s(p))=s(p) [^5] In fact, since the matrix $\mathbf{Z}_{TT}-\frac{\mathbf{x}_{{\mathcal{T}}}\mathbf{x}^{*}_{{\mathcal{T}}}}{4\|\mathbf{x}\|^{2}}$ are Hermitian matrix, all of its eigenvalues are real and non-negative, allowing the inequality hold true. . Taking (C.16) into (C.7) yields
$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\geq$ $~4\left(\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}}-2\frac{\eta}{s(p)}\right)\|\mathbf{x}\|^{2}$
$=$ $~\|\mathbf{x}_{{\mathcal{T}}}\|^{2}-8\frac{\eta}{s(p)}\|\mathbf{x}\|^{2}\overset{(a)}{\geq}(1-16\eta)\|\mathbf{x}_{{\mathcal{T}}}\|^{2},$ (C.17)
where $(a)$ comes from the assumption $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$. Let $\mu_{1}=\sqrt{(1-16\eta)}$, we complete the proof.
Appendix D Proof for Proposition 3
From the definition of $\tilde{f}_{l}$, we can derive that
$|\mathbb{E}[\tilde{f}_{l}]|=\left\{\begin{array}[]{cc}0,&l\notin{\text{supp}(\mathbf{x})},\\ \frac{|\mathbf{x}^{*}\mathbf{z}_{0}||x_{l}|}{4\|\mathbf{x}\|^{2}},&l\in{\text{supp}(\mathbf{x})}.\end{array}\right$. (D.1)
Based on the definition of $S_{\gamma}$, we evaluate the gap between $f_{l},l\in S_{\gamma}$ and $f_{l},l\notin{\text{supp}(\mathbf{x})}$. For $l\in S_{\gamma}$, we have
[…]
$\overset{(a)}{\geq}$ $~\frac{1}{4\|\mathbf{x}\|^{2}}\times\frac{\gamma\|\mathbf{x}\|}{2\sqrt{k}}\times\mu_{1}\sqrt{\frac{1}{2s(p)}}\|\mathbf{x}\|-|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ (D.7)
$\overset{(b)}{>}$ $~\frac{\gamma\mu_{1}}{8\sqrt{2}}\sqrt{\frac{1}{ks(p)}}-\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}}=\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.10)
where $(a)$ is from the definition of $S_{\gamma}$ and Proposition 2 and $(b)$ employs Lemma 6 with $\eta=\frac{\mu_{1}}{8\sqrt{2}}$ and holds with probability exceeding $1-4e^{-c_{\gamma,\mu_{1}}m}$ when $m\geq C_{\mu_{1}}\frac{ks(p)}{\gamma^{2}}\log n$.
On the other hand, for $l\notin{\text{supp}(\mathbf{x})}$, we have
$|f_{l}|\leq|\mathbb{E}[\tilde{f}_{l}]|+|f_{l}+\mathbb{E}[\tilde{f}_{l}]|\overset{(a)}{\leq}\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.11)
where $(a)$ comes from (D.1) and Lemma 6 with $\eta=\frac{\mu_{1}}{8\sqrt{2}}$. By combining (D.10) and (D.11), we claim that all indices in $S_{\gamma}$ are selected in this case. Therefore, we can derive that
$\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$
Appendix E Proof of Proposition 4
Let $\mathbf{z}$ with $\|\mathbf{z}\|=\lambda$ be the eigenvector corresponding to the largest eigenvalue of $\mathbf{Z}_{{\mathcal{T}}}$. Following the same analysis as in Appendix C with the set $\mathcal{T}$} satisfying $|{\mathcal{T}}|\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}$, we obtain similar conclusion to (C.7) and (C.16) as
$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}$ $\geq 4(\tau-\eta_{2})\|\mathbf{z}\|^{2}\|\mathbf{x}\|^{2},$ (E.1)
$\tau$ $\geq\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}}-\eta_{2},$ (E.2)
where we employ $\|\mathbf{z}\|=\lambda$ and Lemma 4 with $g(k,s(p))=1$. Hence (E.1) and (E.2) hold with probability at least $1-2e^{-c_{\gamma,\eta_{2}}m}$ if $m\geq C_{\gamma,\eta_{2}}k\log n$. Suppose $\kappa_{0}=\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}$, taking (E.2) into (E.1) yields
$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\overset{(a)}{\geq}\left(\kappa_{0}-8\eta_{2}\right)(1-\eta_{1})\|\mathbf{x}\|^{4}\overset{\triangle}{=}\mu_{2}^{2}\|\mathbf{x}\|^{4},$ (E.3)
where $(a)$ employs Lemma 3 with $\kappa=1$ and $\beta=\eta_{1}$. In the following, We derive the estimation for $\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})$. From the property of inner product, we have
$\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}=$ $~\|\mathbf{x}_{{\mathcal{T}}}\|^{2}+\|\mathbf{z}\|^{2}-2|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|$
$\leq$ $~(1+\eta_{1}+\kappa_{0})\|\mathbf{x}\|^{2}-2\mu_{2}\|\mathbf{x}\|^{2}.$ (E.4)
Take $\eta_{1}=1-\kappa_{0}$ and $8\eta_{2}=\kappa_{0}-\kappa_{0}^{2}$, (E.4) becomes
$\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}\leq(2-2\sqrt{\kappa_{0}^{3}})\|\mathbf{x}\|^{2}.$ (E.5)
Therefore, for any $0<\delta_{0}<1$, take $\gamma$ as
$1-\frac{\gamma^{2}}{4}=\sqrt[3]{\left(\frac{2-\delta_{0}^{2}}{2}\right)^{2}},$
and the result (E.5) becomes
$\text{dist}(\mathbf{x}_{{\mathcal{T}}},z)\leq\delta_{0}\|\mathbf{x}\|.$ (E.6)
Finally, we have
$\text{dist}(\mathbf{x},\mathbf{z})^{2}$~ $\makebox[29.6192pt][c]{$=$}~\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}+\|\mathbf{x}^{\perp}_{{\mathcal{T}}}\|^{2}$
$\makebox[29.6192pt][c]{$\overset{(a)}{\leq}$}~\delta_{0}^{2}\|\mathbf{x}\|^{2}+\frac{\gamma^{2}}{4}\|\mathbf{x}\|^{2}=\left(\delta_{0}^{2}+\frac{\gamma^{2}}{4}\right)\|\mathbf{x}\|^{2},$
where $(a)$ comes from (E.6) and the definition of $\mathcal{T}$}. Moreover, condition (51) indicates that $\delta\geq\sqrt{\delta_{0}^{2}+\frac{\gamma^{2}}{4}}\in[0,1]$. Therefore, we complete the proof of Proposition 4.

### v3

@@
 Abstract
-Sparse phase retrieval, which aims to recover a $k$-sparse signal from $m$ phaseless measurements, poses a fundamental question regarding the minimal sample complexity required for success.
-While the optimal sample complexity is known to be $\Omega(k\log n)$, existing algorithms can only achieve it for signals under restrictive structural assumptions.
-This paper introduces a novel and robust initialization algorithm, termed generalized Exponential Spectral Pursuit (gESP), designed to overcome this limitation.
-Theoretically, we prove that gESP significantly expands the family of signals that can be recovered with the optimal sample complexity.
-Our analysis unifies existing results on previously studied signal models and surpasses them by establishing performance bounds for a more general class of signals.
-Extensive simulations validate our theoretical findings, demonstrating that gESP consistently outperforms state-of-the-art methods across diverse signal types, thereby pushing the boundaries of efficient and optimal sparse phase retrieval.
-Index Terms:
-Sparse phase retrieval, sample complexity, sparsity, information-theoretical bound.
+Sparse phase retrieval aims to recover a $k$-sparse signal from $m$ phaseless measurements.
+While the theoretically optimal sample complexity for successful recovery is $\Omega(k\log n)$, existing algorithms can only achieve this bound for signals with specific structural assumptions, leading to a notable gap between theory and practice.
+To bridge this gap, we introduce an efficient initialization algorithm, termed generalized Exponential Spectral Pursuit (gESP).
+We prove that gESP can significantly expand the family of signals that are guaranteed to be recovered with the optimal sample complexity, thereby extending the scope of theoretical optimality to a much broader class of signals.
+Extensive simulations validate our theoretical findings and demonstrate that gESP consistently outperforms the state-of-the-art methods across diverse signal types.
 I Introduction
 In many applications, only the intensity information is retained by the measurement system due to hardware limitations [2].
-For example, in optical imaging [2, 3] and X-ray crystallography [4, 5], common sensors such as charge-coupled devices, photosensitive films and human eyes merely record the amplitude of the light wave but ignore the phase.
+For example, in optical imaging [19, 18] and X-ray crystallography [16, 9], common sensors such as charge-coupled devices, photosensitive films, and human eyes merely record the amplitude of the light wave but ignore the phase.
 In these settings, one is faced with the task of recovering the original signal from its magnitude-only measurements, which is often referred to as phase retrieval (PR) [5, 35, 6, 24, 25].
-Succinctly stated, the main task of PR is to reconstruct the $n$-dimensional signal $\mathbf{x}\in\mathbb{C}^{n}$ from a group of phaseless measurements $\mathbf{y}\in\mathbb{R}^{m}$ with
+Succinctly stated, the main task of PR is to reconstruct the $n$-dimensional signal $\mathbf{x}\in\mathbb{C}^{n}$ from a group of phaseless measurements $\mathbf{y}\in\mathbb{R}^{m}$ given by
 $y_{i}=|\mathbf{a}_{i}^{*}\mathbf{x}|,\ i=1,2,\cdots m,$ (1)
 where $\mathbf{a}_{i}\in\mathbb{C}^{n}$ is the measurement vector.
-In this paper, we consider the complex Gaussian measurements, i.e., $\mathbf{a}_{i}\sim\mathcal{CN}(n)$, and $\mathbf{x}$ is $k$-sparse, which means that there are at most $k$ non-zero entries in $\mathbf{x}$.
-This problem is the so-called sparse phase retrieval problem.
-Researchers are interested in the minimum measurement number that one needs to reconstruct the signal [11, 12, 13, 14].
-It has been shown that with m=4k2m=4k-2 phaseless measurements [15], the non-linear system (1) with a $k$-sparse prior produces the unique solution based on generic sensing measurements.
-Moreover, researches have proved that
+In this paper, we consider complex Gaussian measurements, i.e., $\mathbf{a}_{i}\sim\mathcal{CN}(n)$, and $\mathbf{x}$ is $k$-sparse, implying that there are at most $k$ non-zero entries in $\mathbf{x}$.
+This problem is known as the sparse phase retrieval problem.
+Researchers are interested in the minimum number of measurements required to reconstruct the signal [30, 31, 10, 17].
+It has been shown that with m=4k2m=4k-2 phaseless measurements [28], the non-linear system (1) with a $k$-sparse prior produces a unique solution based on generic sensing measurements.
+Moreover, it has been proven that
 $m=\Omega(k\log n)$ (2)
 complex Gaussian measurements are sufficient to guarantee the stability of the solution [8, 22]. This sample complexity is often called the information-theoretical bound for sparse phase retrieval.
-For polynomial algorithms, they appear to require more measurements for the exact recovery.
-Back in 2013, Li and Voroninski [18] utilized the matrix lifting technique and convex relaxation methods to solve the sparse phase retrieval problem.
+For polynomial-time algorithms performing exact recovery of $\mathbf{x}$, it appears to require significantly more phaseless measurements.
+Back in 2013, Li and Voroninski [14] utilized a matrix lifting technique and convex relaxation methods to solve the sparse phase retrieval problem.
 They proved that it requires
@@
 While their sample complexity revealed a clear discrepancy compared to (2), the authors were unsure whether this difference was inherent to the problem itself or simply a consequence of limitations in their modeling approach.
-This is afterwards called the statistical-to-computational gap.
-This gap is further investigated in the thresholded Wirtinger Flow (TWF) algorithm [19], which is a typical greedy algorithm containing two stages — the initialization stage and the refinement stage.
-They showed that both stages require $m=\Omega(k^{2}\log n)$ samples for the exact final recovery.
-This two-stage approach was followed in subsequent algorithms, e.g., SPARse Truncated Amplitude flow (SPARTA[20], Compressive Phase Retrieval with Alternating Minimization (CoPRAM[21], and Subspace Phase Retrieval (SPR[22].
-It is shown via some geometric analysis [23, 22, 24] and sophisticated optimization approaches [25] that the sample complexity of the refinement stage can be improved to the information-theoretical bound (2), if a $\delta$-neighborhood estimate of the target signal is provided.
+Later, this discrepancy was identified as the statistical-to-computational gap [22, 21], particularly in the context of generic Gaussian measurements.
+Nevertheless, it has been shown that this barrier can be circumvented by imposing specific structures on the measurement process.
+Prior works, such as [1, 11], demonstrated that the optimal sample complexity is achievable in polynomial time when additional constraints are placed on the sensing vectors.
+However, under the standard unstructured Gaussian setting, verifying the existence of this gap remains a significant challenge.
+This gap was further investigated in the thresholded Wirtinger Flow (TWF) algorithm [4], which is a typical greedy algorithm containing two stages: the initialization stage and the refinement stage.
+The authors showed that both stages require $m=\Omega(k^{2}\log n)$ samples for exact final recovery.
+This two-stage approach was followed in subsequent algorithms, e.g., SPARse Truncated Amplitude flow (SPARTA) [27], Compressive Phase Retrieval with Alternating Minimization (CoPRAM) [12], and Subspace Phase Retrieval (SPR) [32].
+It is shown via some geometric analysis [21, 32, 15] and sophisticated optimization approaches [20] that the sample complexity of the refinement stage can be improved to the information-theoretical bound (2), provided that a $\delta$-neighborhood estimate of the target signal is available.
 However, to achieve this, the initialization stage requires
@@
 More precisely, it is the requirement of finding an estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ that leads to the statistical-to-computational gap.
-Despite the persistence of this gap, numerous efforts are underway to achieve improved results. For the sake of notational simplicity, we define $s(p)$ as
+Despite the persistence of this gap, numerous efforts are underway to achieve improved results. For the sake of notational simplicity, we define the structural function $s(p)$ as
 $s(p)=\frac{\|\mathbf{x}\|^{2}}{\sum_{j=1}^{p}|x_{(j)}|^{2}},~~p=1,2,\cdots,n,$ (5)
-where $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ are the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module.
+where $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ are the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes.
 A formal definition and the corresponding interpretation are presented in Section III-A.
-Wu and Rebeschini [27] presented the algorithm call Hadamard Wirtingle Flow (HWF).
-They showed that the gap vanished for some specific structure of the target signal $\mathbf{x}$.
+Wu and Rebeschini [29] presented an algorithm called Hadamard Wirtinger Flow (HWF).
+They showed that the gap vanished for signals with specific structures.
 Specifically, by assuming that the minimum magnitude of non-zeros $x_{(k)}$ is on the order of $\frac{\|\mathbf{x}\|}{\sqrt{k}}$, i.e.,
 $\left\lvert x_{(k)}\right\rvert=\Omega\left(\frac{\|\mathbf{x}\|}{\sqrt{k}}\right),$ (6)
-they show that the sample complexity can be reduced to
+they showed that the sample complexity can be reduced to
 $m=\Omega\left(\max\left\{ks(1)\log n,\sqrt{ks(1)}\log^{3}n\right\}\right).$ (7)
 This means that when the maximum entry $x_{(1)}$ of $\mathbf{x}$ is on the order of $\|\mathbf{x}\|$ (ignoring the $\log^{3}n$ term), (7) can be further reduced to (2), and thus reach the theoretical lower bound.
-This result (7) is further improved by Cai et al.
+The result of (7) was further improved by Cai et al.
 [3] and Xu et al.
@@
 In particular, Cai et al.
-[26] proposed Truncated Power method (TPM) to remove the requirements on the minimum magnitude of non-zeros (i.e., (6)).
+[3] proposed the Truncated Power Method (TPM) to remove the requirements on the minimum magnitude of non-zeros (i.e., Eq.
+(6)).
 Xu et al.
-[28] introduce an exponential spectrum and proposed Exponential Spectral Pursuit (ESP) algorithm to eliminate the $\log^{3}n$ term in (7).
-Finally, the sample complexity is reduced to
-$m=\Omega\left(ks(1)\log n\right)$ (8)
-samples are sufficient for ESP to produce a good estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ without any assumption on $x_{(k)}$.
+[33] introduced an exponential spectrum and proposed the Exponential Spectral Pursuit (ESP) algorithm to eliminate the $\log^{3}n$ term in (7).
+The sample complexity was reduced to
+$m=\Omega\left(ks(1)\log n\right),$ (8)
+which is sufficient for ESP to produce a good estimate falling into the $\delta$-neighborhood of $\mathbf{x}$ without any assumption on $x_{(k)}$.
 However, this result only achieves the optimal bound (2) for a restrictive class of signals, i.e., those with a single dominant entry where $s(1)=\Theta(1)$.
@@
 In this paper, we develop the generalized Exponential Spectral Pursuit (gESP) algorithm to address this challenge.
-Specifically, gESP selects an index set rather than a single index in its first step (i.e., Step 1 in Table 1).
+Specifically, gESP selects an index set rather than a single index in its first step (i.e., Step 1 in Algorithm 1).
 This key modification allows gESP to achieve the optimal sample complexity for a much broader class of signals.
@@
 $m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$ (10)
-It can be shown that in the case where $s(1)=\Theta(\sqrt{k})$ and $s(\sqrt{k})=\Theta(1)$, this result will also be better than (8). The detailed discussion can be found in Section IV.
+It can be shown that in the case where $s(1)=\Theta(\sqrt{k})$ and $s(\sqrt{k})=\Theta(1)$, this result will also be better than (8). A detailed discussion can be found in Section IV.
 The remainder of this paper is organized as follows.
@@
 Section III contains theoretical results for the sample complexity of gESP in different cases and makes simple comparisons between these results.
-In Section IV, discussions are directed in several aspects to illustrate the superiority of our result.
+In Section IV, we discuss various aspects to illustrate the superiority of our result.
 In Section V, the proofs of the main theorems are given based on the analysis for each step of gESP.
@@
 Let $[n]$ denote the set $\{1,2,3,\cdots n\}$.
-We denote $x_{(i)}$ as the $i$-largest entry of $\mathbf{x}$ in magnitude.
-For the sake of readability, we denote $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ as the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module.
+For the sake of readability, we denote $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ as the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes.
 For any set $\subset[n], S{\mathbb{C}}^{S}$ is defined as the subspace of ${\mathbb{C}}^{n}$ spanned by vectors supported on $S$, i.e., $\{\mathbf{x}|\mathbf{x}\in{\mathbb{C}}^{n},{\text{supp}(\mathbf{x})}=S\}$.
@@
 For the matrix $\mathbf{A}$, we define $\mathbf{A}_{S}$ as the matrix which keeps columns and rows indexed by $S$ while setting others to zero.
-Given a vector $\mathbf{x}$, the conjugate transpose, $\ell_{2}$ norm and $\ell_{0}$ norm of $\mathbf{x}$ are denoted as $\mathbf{x}^{*}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{0}$ respectively.
+Given a vector $\mathbf{x}$, the conjugate transpose, $\ell_{2}$ norm, and $\ell_{0}$ norm of $\mathbf{x}$ are denoted as $\mathbf{x}^{*}$, $\|\mathbf{x}\|$ and $\|\mathbf{x}\|_{0}$ respectively.
 An $n$-dimensional standard complex Gaussian random vector, denoted as $\mathbf{a}\sim\mathcal{CN}(n)$, is defined as $\mathbf{a}=\mathbf{a}_{1}+\mathrm{j}\mathbf{a}_{2}$, with $\mathbf{a}_{1},\mathbf{a}_{2}\sim\mathcal{N}(\mathbf{0},\frac{1}{2}\mathbf{I})$, and j is the imaginary unit.
 We use $a\sim b$ to represent that $a$ differs from $b$ by a constant factor.
-Following the convention in the literature, we use standard asymptotic notations to characterize the complexity as bellow.
+Throughout this paper, $c$ and $C$ denote universal positive constants whose values may change from line to line.
+We say an event occurs with high probability if it holds with probability at least $1-n^{-c}$ for some constant $c>0$.
+Following the convention in the literature, we use standard asymptotic notations to characterize the complexity as follows:
   • Big-Omega Notation ($\Omega$): We say $f(n)=\Omega(g(n))$ if there exist positive constants $c$ and $n_{0}$ such that for all $n\geq n_{0}$, we have $0\leq c\cdot g(n)\leq f(n)$.
@@
 $\left\lvert\mathbf{x}_{(k)}\right\rvert=\Omega\left(\frac{\|\mathbf{x}\|}{\sqrt{k}}\right).$ (11)
-This assumption is made for the convenience of our brief introduction. One can find from our theorems that this assumption is actually unnecessary; also see the results in [26, 28].
+While this assumption is made for the convenience of our brief introduction, it will be shown in our main theorems that this assumption is actually unnecessary; Similar arguments can be found in [3, 33].
 II-A Prior work
@@
   1. 1.
-    Earlier researches contain two steps for initialization.
+    Earlier initialization approaches typically involve two steps:
     Step 1: Construct a spectrum $\mathbf{Z}$ and take the diagonal entries to be the set $\{Z_{j}\}_{j=1}^{n}$.
 Sorting $\{Z_{j}\}_{j=1}^{n}$ and taking the first $k$ indices yields the final support estimate $S$.
-    Generally, the expectation of $\mathbf{Z}$ takes the form as follow
+    Generally, the expectation of $\mathbf{Z}$ takes the following form:
     $\mathbb{E}[\mathbf{Z}]=\alpha\|\mathbf{x}\|^{2}\mathbf{I}+\beta\mathbf{x}\mathbf{x}^{*}.$ (12)
@@
     where $\alpha,\beta$ are both numerical constants, determined by the specific definition of $Z_{j}$.
-For instance, [20, 29] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}|\mathbf{a}_{ij}|^{2}$ and the corresponding $\alpha=1,\beta=2$, whereas [14] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}|\mathbf{a}_{ij}|$ and the corresponding $\alpha=\frac{2}{\pi\|\mathbf{x}\|},\beta\approx\frac{1}{6\|\mathbf{x}\|}$.
+For instance, [27, 34] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}|\mathbf{a}_{ij}|^{2}$ with the corresponding $\alpha=1,\beta=2$, whereas [17] set $Z_{j}=\frac{1}{m}\sum_{i=1}^{m}y_{i}|\mathbf{a}_{ij}|$ with the corresponding $\alpha=\frac{2}{\pi\|\mathbf{x}\|},\beta\approx\frac{1}{6\|\mathbf{x}\|}$.
 It can be seen that there exists an apparent gap between $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})}$ and $\{\mathbb{E}[Z_{j}]\}_{j\in\operatorname*{supp}(\mathbf{x})^{c}}$.
 Specifically,
-    $\min_{j\in\operatorname*{supp}(\mathbf{x})}\mathbb{E}[Z_{j}]-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}\mathbb{E}[Z_{j}]=$ $\beta|{x_{(k)}}|^{2}$
-    $\sim$ $\frac{\|\mathbf{x}\|^{2}}{k}.$ (14)
+    $\min_{j\in\operatorname*{supp}(\mathbf{x})}\mathbb{E}[Z_{j}]-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}\mathbb{E}[Z_{j}]=$ $~\beta|{x_{(k)}}|^{2}$
+    $\sim$ $~\frac{\|\mathbf{x}\|^{2}}{k}.$ (14)
     Thus, employing some concentration inequalities on $Z_{j}$ and $\mathbb{E}[Z_{j}]$ finally leads to
@@
     which means one can obtain the estimate of support set by sorting the elements of $\{Z_{j}\}_{j=1}^{n}$ and selecting the indices corresponding to the $k$ largest values.
-Then it can be guaranteed that this support estimate contains all the true candidates.
-Theoretical analysis shows that deriving an appropriate concentration from the gap $\frac{\|\mathbf{x}\|^{2}}{k}$ (see (14)) requires the sample complexity being [^1] When $Z_{j}$ is statistically heavy-tailed ((e.g., involving the forth power of Gaussian), some extra term containing $\log^{3}n$ (e.g., see (7)) will appear in the complexity.
+This guarantees that the support estimate contains all the true candidates.
+Theoretical analysis shows that deriving an appropriate concentration from the gap $\frac{\|\mathbf{x}\|^{2}}{k}$ (see (14)) requires the sample complexity being [^1] When $Z_{j}$ is statistically heavy-tailed (e.g., involving the fourth power of Gaussian), some extra term containing $\log^{3}n$ (e.g., see (7)) will appear in the complexity.
 This is because one needs to use a truncation argument to analyze the tail bound.
@@
     which results from the re-scaled gap $\frac{|{x_{(k)}}|^{2}}{\|\mathbf{x}\|^{2}}\leq\frac{1}{k}$.
-    Step 2: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
-    The maximal eigenvector of $\mathbb{E}[\mathbf{Z}]$ is $\mathbf{x}$ and $\mathbb{E}[\|\mathbf{z}\|^{2}]=\|\mathbf{x}\|^{2}$.
+    Step 2: Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
+    The leading eigenvector of $\mathbb{E}[\mathbf{Z}]$ is $\mathbf{x}$ and $\mathbb{E}[\|\mathbf{z}\|^{2}]=\|\mathbf{x}\|^{2}$.
 Therefore, we shall obtain a good estimate of $\mathbf{x}$.
@@
   2. 2.
-    Recently, a novel initialization paradigm is proposed and developed in [27, 26, 28], which contains four steps.
-Instead of directly estimating the support in one step, an index $j_{\text{max}}$ will be firstly selected, which corresponds to the largest value among $Z_{j}$’s.
+    Recently, a novel initialization paradigm was developed in [29, 3, 33], which contains four steps.
+Instead of directly estimating the support in one step, an index $j_{\text{max}}$ is first selected corresponding to the largest value among $Z_{j}$’s.
     Step 1: Construct a spectrum $\mathbf{Z}$ and take the diagonal entries to be the set $\{Z_{j}\}_{j=1}^{n}$. Select the index $j_{\text{max}}$ corresponding to the largest value among $Z_{j}$’s.
@@
     $=\alpha\|\mathbf{x}\|^{2}\mathbf{e}_{j_{\text{max}}}+\beta{x_{(1)}}\mathbf{x}.$ (20)
-    Establishing some concentration inequalities on this random variable yields the similar result to (15)
+    Establishing some concentration inequalities on this random variable yields a result similar to (15):
     $\min_{j\in\operatorname*{supp}(\mathbf{x})}|\mathbf{Z}\mathbf{e}_{j_{\text{max}}}|_{j}-\max_{j\in\operatorname*{supp}(\mathbf{x})^{c}}|\mathbf{Z}\mathbf{e}_{j_{\text{max}}}|_{j}\approx\$$\beta|{x_{(1)}}{x_{(k)}}|$
     $\sim\$$\frac{|{x_{(1)}}|\|\mathbf{x}\|}{\sqrt{k}}.$ (21)
-    To identify the true support set for (21), it requires
+    To identify the true support set based on (21), it requires
     $m=\Omega\left(k\frac{\|\mathbf{x}\|^{2}}{|{x_{(1)}}|^{2}}\log n\right).$ (22)
-    Step 4: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
-    This step is as the same as Step 2 of the previous method.
+    Step 4: Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
+    This step is the same as Step 2 of the previous method.
     Finally, the overall sample complexity of this method is
@@
 Note that the term $|{x_{(1)}}|$ in this gap arises from the fact that we simply select one maximal index in Step 1.
-If more significant indices are chosen in Step 1, the final gap will indeed be enlarged, and thus decrease the complexity in Step 3333In fact, the complexity in Step 1 will increase while that in Step 3 will decrease as more significant indices are chosen.
-This means that there is a trade-off in the number of the indices that should be chosen in Step 1..
-Then the remaining issue is to design Step 2.
-Assume that we obtain the set $S^{0}$ from Step 1.
+If more significant indices are chosen in Step 1, the final gap will indeed be enlarged, thereby decreasing the complexity in Step 3333In fact, the complexity in Step 1 will increase while that in Step 3 will decrease as more significant indices are chosen.
+This means that there exists a trade-off in the number of the indices that should be chosen in Step 1..
+Then the remaining issue is to design Step 2, where the set $S^{0}$ from Step 1 may contain more than one support candidate.
 Following the previous method, a straightforward idea is to construct the vector $\mathbf{e}^{0}$ such that it takes $1$ on $S^{0}$ and 0 for the rest.
-A more sophisticated design, however, is to compute the maximal eigenvector of $\mathbf{Z}_{S^{0}}$ as $\mathbf{e}^{0}$.
-This also matches the previous method since $\mathbf{e}_{j_{\text{max}}}$ is exactly the maximal eigenvector of $\mathbf{Z}_{\{j_{\text{max}}\}}$.
-So far, we have completed the introduction of the modification idea.
+A more sophisticated design, however, is to compute the leading eigenvector of $\mathbf{Z}_{S^{0}}$ as $\mathbf{e}^{0}$.
+This also matches the previous method since $\mathbf{e}_{j_{\text{max}}}$ is exactly the leading eigenvector of $\mathbf{Z}_{\{j_{\text{max}}\}}$.
+So far, we have completed the introduction of the proposed modification.
 The specific algorithm is summarized in Algorithm 1.
@@
 The more detailed sample complexity and numerical conclusions are formally presented in Section III.
-The spectrum in our proposed gESP algorithm is
+Following [33], we utilize the spectrum:
 $\mathbf{Z}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right)\mathbf{a}_{i}\mathbf{a}_{i}^{*},$ (24)
@@
 $\|\mathbf{x}_{S^{0}}\|^{2}\sim\sum_{i=1}^{p}|x_{(i)}|^{2}.$ (27)
-Since the maximal eigenvector of $\mathbb{E}[\mathbf{Z}_{S^{0}}]$ is $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$, we can approximate $\mathbf{e}_{0}$ as $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$.
+Since the leading eigenvector of $\mathbb{E}[\mathbf{Z}_{S^{0}}]$ is $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$, we can approximate $\mathbf{e}_{0}$ as $\mathbf{x}_{S^{0}}/\|\mathbf{x}_{S^{0}}\|$.
 Then, Step 2 ensures that
@@
 $\approx$~ $\beta|\mathbf{x}^{*}\mathbf{e}^{0}|\cdot{x_{(k)}}\sim\frac{\sqrt{\sum_{i=1}^{p}|x_{(i)}|^{2}}\|\mathbf{x}\|}{\sqrt{k}}.$ (30)
-Therefore, Step 3 along with Step 4 ensure that $\mathbf{z}$ is a good initialization of the target signal $\mathbf{x}$.
-Algorithm 1  Generalized Exponential Matching Pursuit
-Input: $\mathbf{A}$, $\mathbf{y}$, the sparsity $k$, the parameter pp.
+This means that in Step 3, we can distinguish the true support entries. Then Step 4 can ensure that $\mathbf{z}$ is a good initialization of the target signal $\mathbf{x}$.
+Algorithm 1  Generalized Exponential Spectral Pursuit
+Input: $\mathbf{A}$, $\mathbf{y}$, the sparsity $k$, the parameter $p$.
 Step 1:
 Generate an index set $S^{0}$ corresponding to the largest $p$ diagonal elements of $\mathbf{Z}$.
-Step 2: Set $\mathbf{e}^{0}$ as the unit maximal eigenvector of $\mathbf{Z}_{S^{0}}$.
+Step 2: Set $\mathbf{e}^{0}$ as the unit leading eigenvector of $\mathbf{Z}_{S^{0}}$.
 Step 3:
@@
 Step 4:
-Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
+Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
 Output: $\mathbf{z}$.
@@
 where $\mathrm{j}$ is the imaginary unit.
-In the initialization stage, typical algorithms such as HWF, SPARTA and TP usually aim at producing a good estimate $\mathbf{z}$ falling into the $\delta$ neighborhood of the target signal, i.e.,
+In the initialization stage, typical algorithms, such as HWF, SPARTA and TPM, usually aim to produce a good estimate $\mathbf{z}$ falling into the $\delta$-neighborhood of the target signal, i.e.,
 $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|,$ (32)
-which is the prerequisite of the refinement stage [19, 20, 29]. In the analysis of gESP, we follow the same line and focus on obtaining the estimate satisfying (32).
+which is the prerequisite of the refinement stage [4, 27, 34]. In the analysis of gESP, we follow the same line and focus on obtaining an estimate satisfying (32).
 Furthermore, we give the definition of the structure function $s(p)$ of the target signal $\mathbf{x}$.
 Definition 2.
-Let $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ be the rearrangement of the entries of $\mathbf{x}$ with a descend order in their module. the structure function $s(p)$ is given by
+Let $x_{(1)},x_{(2)},\cdots x_{(k)},\cdots x_{(n)}$ be the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes. The structure function $s(p)$ is given by
 $s(p)=\frac{\|\mathbf{x}\|^{2}}{\sum_{j=1}^{p}|x_{(j)}|^{2}},~~p=1,2,\cdots,n.$ (33)
 The structure function, $s(p)$, quantifies the structural information of a signal $\mathbf{x}$ by measuring the energy distribution across its components.
-Specifically, $s(p)$ represents the proportion of the total signal energy captured by the $p$ largest magnitude components.
+Specifically, $s(p)$ represents the inverse of the proportion of the total signal energy captured by the $p$ largest magnitude components.
 Therefore, the behavior of $s(p)$ as $p$ varies directly reflects how energy is concentrated or dispersed within the signal, thus characterizing its structure.
@@
   1. 1.
-    $s(p)$ is monotonously decreasing with respect to $p$. Furthermore,
+    $s(p)$ is monotonically decreasing with respect to $p$. Furthermore,
     $1\leq s(p)\leq\frac{k}{p}.$ (34)
@@
     $ps(p)=\frac{\|\mathbf{x}\|^{2}}{\frac{1}{p}\sum_{i=1}^{p}|x_{(i)}|^{2}},$ (35)
-    which implies $ps(p)$ is monotonously increasing with respect to $p$. Furthermore,
+    which implies $ps(p)$ is monotonically increasing with respect to $p$. Furthermore,
     $p\leq ps(p)\leq k.$ (36)
+Finally, to facilitate rigorous theoretical analysis and decouple statistical dependencies among different iterations, we employ a standard sample splitting strategy [13, 5] in our proof.
+Specifically, the measurement set $\mathcal{D}=\{(a_{i},y_{i})\}_{i=1}^{m}$ is randomly partitioned into $4$ disjoint subsets $\mathcal{D}_{1},\dots,\mathcal{D}_{4}$, where each subset is used for a single step of our algorithm.
+Here we emphasize that sample splitting is purely a theoretical artifact used to simplify the probabilistic analysis.
+In our practical implementation (Algorithm 1) and numerical experiments, we utilize the full dataset $\mathcal{D}$ at every step to maximize the sample efficiency.
+Since the number of steps is small and fixed (i.e., $4$), splitting the data only affects the sample complexity by a constant factor, leaving the order of the convergence rate unchanged.
 III-B Theoretical results for sample complexity of gESP
@@
 Theorem 1.
-For any numerical constant $0<\delta<1$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ with probability at least $1-4e^{-c_{\delta}m}$.
+For any constant $0<\delta<1$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ with probability at least $1-n^{-c}$.
 The corresponding required sample complexity depends on an input parameter $p\in[k]$; for a general $p$, it is given by
@@
 Remark 1.
-The state-of-the-art result [28] is
+The state-of-the-art result [33] is the special case $p=1$ of our result, which is
 $m$ $=\Omega\left(\max\left\{s^{2}(1)\log n,ks(1)\log n\right\}\right)$
@@
 Therefore, our result (39) exhibits a significant improvement.
+Remark 2.
+The statistical-to-computational gap persists for signals with maximally delocalized energy.
+Specifically, the worst-case sample complexity of $\Omega(k^{2}\log n)$ is required if and only if the signal is almost “flat”, i.e., $s(p)=\Theta\left(\frac{k}{p}\right)$ for all $p\in[1,k]$.
 Remark 3.
@@
 Since $s(p)\geq 1$ (see (34)), (41) implies $s(p)=\Theta(1)$ for some $p\in[k]$.
-Then, (42) requires $p\leq c\lceil\sqrt{k}\rceil$ for some constant $c$.
-Moreover, if $s(p)=\Theta(1)$ holds for some fixed $p\leq c\lceil\sqrt{k}\rceil$, then $s(\lceil\sqrt{k}\rceil)=\Theta(1)$ since $s(p)$ is monotonously decreasing and larger than $1$.
-Therefore, to achieve the information-theoretical bound, gESP requires $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
+Then, (42) requires $p\leq\lceil c\sqrt{k}\rceil$ for some constant $c$.
+Moreover, if $s(p)=\Theta(1)$ holds for some fixed $p\leq\lceil c\sqrt{k}\rceil$, then $s(\lceil\sqrt{k}\rceil)=\Theta(1)$ since $s(p)$ is monotonically decreasing and larger than $1$.
+Therefore, to achieve the information-theoretical bound (2), gESP requires $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
 This result provides the general sample complexity for gESP.
-However, in the case where $s(p)$ is unknown, determining the optimal $p$ is challenging.
-One compromise is to execute Algorithm 1 for $k$ times with $p$ taking each value in $[k]$ in turn.
-Then we shall obtain a set containing $k$ estimates of the target signal $\mathbf{x}$.
-It can be guaranteed that at least one estimate falls into the $\delta$-neighborhood of $\mathbf{x}$ with the same sample complexity requirement as (39).
-The formal statement is in the following corollary.
+Interestingly, this result depends solely on the cumulative energy profile $s(p)$, imposing no lower bound requirement on the minimum non-zero magnitude $|x_{(k)}|$.
+This is due to that the assumption (11) is required for exact support recovery, whereas our analysis only targets an $\ell_{2}$-distance bound ($\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$), which, therefore, can be met under a relaxed requirement.
+A more specific explanation is provided below Proposition 3.
+In the case where $s(p)$ is unknown, determining the optimal $p$ is challenging.
+One compromise is to execute Algorithm 1 $k$ times with $p$ taking each value in $[k]$.
+This yields a set containing $k$ estimates of the target signal $\mathbf{x}$.
+Naturally, at least one estimate falls into the $\delta$-neighborhood of $\mathbf{x}$ with the same sample complexity requirement as (39), which leads to the following corollary.
 Corollary 1.
-Execute Algorithm 1 for $k$ times with $p$ taking each value in $[k]$ in turn, and collect all the estimates into a set $\mathcal{X}$.
-Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, there exists at least one estimate $\mathbf{z}\in\mathcal{X}$ that satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ when
+Run Algorithm 1 $k$ times with $p$ taking each value in $[k]$ in turn, and collect all the estimates into a set $\mathcal{X}$.
+Then, for any $0<\delta<1$, with probability at least $1-n^{-c}$, there exists at least one estimate $\mathbf{z}\in\mathcal{X}$ that satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ when
 $m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p)\log n,ks(p)\log n\right\}\right).$ (43)
-When multiple executions of the algorithm with varying $p$ are impractical due to computational constraints, we are compelled to explore the sample complexity for specific, analytically tractable choices of $p$.
+Remark 4.
+While Corollary 1 guarantees the optimal sample complexity by iterating $p$ over the entire range of $[1,k]$, this approach entails running the algorithm $k$ times.
+To mitigate the computational cost in practice, one can adopt a logarithmic search strategy, namely, selecting candidates from a subset $\mathcal{P}=\{1,2,4,\dots,2^{\lfloor\log_{2}k\rfloor}\}$.
+Since the complexity bound varies smoothly with $p$, this heuristic typically identifies a parameter that yields a sample complexity within a constant factor of the global optimum, while significantly reducing the computational overhead from $O(k)$ to $O(\log k)$ rounds.
+Due to computational constraints, multiple executions of the algorithm with varying $p$ are impractical.
+Thus, we are motivated to explore the sample complexity for specific, analytically tractable choices of $p$.
 A natural starting point for such an analysis is the intuitive choice of setting $p=k$.
@@
 Take $p=k$ as the input of Algorithm 1.
-Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number of samples $m$ satisfies
+Then, for any $0<\delta<1$, with probability at least $1-n^{-c}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number of samples $m$ satisfies
 $m=\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right).$ (44)
@@
 Take $p=\lceil\sqrt{k}\rceil$ as the input of Algorithm 1.
-Then, for any $0<\delta<1$, with probability at least $1-4e^{-c_{\delta}m}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number $m$ of samples satisfies
+Then, for any $0<\delta<1$, with probability at least $1-n^{-c}$, the output $\mathbf{z}$ of Algorithm 1 satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ provided that the number $m$ of samples satisfies
 $m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$ (45)
-One one hand, the result (45) cannot be worse than (44) since
+On one hand, the result (45) cannot be worse than (44) since
 $~\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right)$
@@
 $=$ $~\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right).$ (46)
-Moreover, there exist some signals that make the less-than sign strictly holds. Therefore, (45) is indeed tighter than (44). See Section IV-B for an in-depth discussion.
+Moreover, there exist some signals that make the inequality strictly hold. Therefore, (45) is indeed tighter than (44). See Section IV-B for an in-depth discussion.
 On the other hand, compared to the state of the art, the results (44) and (45) are superior for certain signals.
-In other words, even though $s(p)$ is unknown as well as executing Algorithm 1 for multiple times is infeasible, our algorithm taking $p=k$ or $\lceil\sqrt{k}\rceil$ can perform better compared with existing algorithms.
-Discussion IV-A gives an in-depth discussion.
+In other words, even though $s(p)$ is unknown and executing Algorithm 1 multiple times is infeasible, our algorithm taking $p=k$ or $\lceil\sqrt{k}\rceil$ can perform better compared with existing algorithms.
+Section IV-A provides an in-depth discussion.
 IV Discussion
 IV-A Broader Conditions for Optimal Sample Complexity
-As has been discussed in Remark 1 and Remark 2, gESP makes a significant improvement over the state-of-the-art result, if the structural function $s(p)$ is known.
+As discussed in Remark 1 and Remark 3, gESP offers a significant improvement over the state-of-the-art result, provided that the structural function $s(p)$ is known.
 This subsection further explores this superiority, demonstrating that gESP achieves the optimal sample complexity for a broader class of signals.
@@
 $m=\Omega\left(ks(1)\log n\right)$ (47)
-requires $s(1)=\Theta(1)$, which is a more strict condition.
-The improvement results from taking $p$ entries in Step 1 into consideration, thus relaxing the condition on $x_{(1)}$ to $x_{(1)},x_{(1)},\dots,x_{(p)}$.
+requires $s(1)=\Theta(1)$, which is a stricter condition.
+The improvement results from considering $p$ entries in Step 1, thus relaxing the condition on $x_{(1)}$ to the set $\{x_{(1)},\dots,x_{(p)}\}$.
 Fig.
-1 gives a comprehensive illustration.
+1 provides a comprehensive illustration.
 Figure 1: An illustration of the relaxed condition for achieving the optimal sample complexity $\Omega(k\log n)$.
 While state-of-the-art methods require the signal’s energy to be concentrated in its single largest entry: $s(1)=\Theta(1)$, gESP achieves optimality for a broader class of signals where the energy is concentrated among the top $\lceil\sqrt{k}\rceil$ entries: $s(\lceil\sqrt{k}\rceil)=\Theta(1)$.
-Moreover, even in the case where $s(p)$ is unknown, our results (Corollary 2 and 3) can also perform better in many cases.
+Moreover, even when $s(p)$ is unknown, our results (Corollary 2 and 3) can still perform better in many cases.
 In the following, we set $p=k$ and consider Corollary 2 for a simple discussion, although we can prove the strict superiority of the result derived in Corollary 3 (see Section IV-B).
@@
 $m=\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right),$ (48)
-Therefore, (48) outperforms the state-of-the-art result (47) once $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$. Here, we offer an example to demonstrate the existence of such signals.
+Consequently, (48) outperforms the state-of-the-art result (47) once $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$. To illustrate this, we construct a parametric family of signals as follows.
 Example 1.
-Assume that both $\sqrt{k}$ and $\sqrt[6]{k}$ are integers [^4] In fact, we can also construction such examples even though this assumption does not hold.
-However, omitting it will lead to unnecessary complications in the construction (e.g., involving terms like $\lceil\sqrt{k}\rceil\pm 1$).
-The same phenomenon also occurs in Example 2.
-For the sake of simplicity, we make this assumption.
-The target signal $\mathbf{x}$ satisfies
-$|x_{(i)}|^{2}=\begin{cases}\frac{\|\mathbf{x}\|^{2}}{\sqrt{k}},&i=1,\\ \frac{\|\mathbf{x}\|^{2}}{\sqrt{k}-1}\times\left(\frac{1}{\sqrt[6]{k}}-\frac{1}{\sqrt{k}}\right),&i=2,\cdots,\sqrt{k},\\ \frac{\|\mathbf{x}\|^{2}}{k-\sqrt{k}}\times\left(1-\frac{1}{\sqrt[6]{k}}\right),&i=\sqrt{k}+1,\cdots,k.\end{cases}$
-In this case, $s(1)=\sqrt{k}$ and $s(\sqrt{k})=\sqrt[6]{k}$, and one can thus find (48) outperforms (47).
+Consider a family of $k$-sparse signals $\mathbf{x}\in\mathbb{R}^{n}$ parameterized by a decay exponent $\delta\in(1/4,1/2)$. Let
+$x_{(i)}=\begin{cases}k^{-\delta},&i=1,\\[1.99997pt] \frac{1}{2}k^{-\delta},&2\leq i\leq\lceil\sqrt{k}\rceil,\\[1.99997pt] k^{-1/2},&\lceil\sqrt{k}\rceil<i\leq k,\\[1.99997pt] 0,&i>k.\end{cases}$ (49)
+For this family of signals, it holds that $s(1)=\Theta(k^{2\delta})$ whereas $s^{2}(\lceil\sqrt{k}\rceil)=\Theta(k^{4\delta-1})$, and thus we have $\Theta(s^{2}(\lceil{\sqrt{k}}\rceil))<\Theta(s(1))$ with the improvement factor being $\Theta(k^{1-2\delta})$.
+This family represents signals where most of the signal’s energy is spread fairly evenly across a long “flat tail” of many small nonzero entries, while only a small fraction sits in the first $\sqrt{k}$ larger entries, so it looks like a modest spike plus a short plateau on top of a broad, energy-dominating background.
+The proof for this example is provided in Appendix A-C.
 Interestingly, when the condition $s(\lceil{\sqrt{k}}\rceil)=\Theta(1)$ holds, our result (48) also achieves the information-theoretic bound (2).
-Significantly, even without prior knowledge of the signal structure $s(p)$, this optimality is attainable for the same class of signals in the case $s(p)$ is known, simply by setting the parameter $p=k$.
+Notably, even without prior knowledge of the signal structure $s(p)$, this optimality is attainable for the same class of signals as in the case where $s(p)$ is known, simply by setting the parameter $p=k$.
 IV-B Strict Superiority of Corollary 3 over Corollary 2
@@
 $m=\Omega\left(\min_{p\in[\lceil\sqrt{k}\rceil]}\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right),$
-is better than (48) in Section III-B. In this section, we provide one certain example to illustrate this strict superiority.
+is better than (48) in Section III-B. In this section, we provide a concrete example to illustrate this strict superiority.
 Example 2.
-Assume that both $\sqrt{k}$ and $\sqrt[4]{k}$ are integers. Construct the target signal $\mathbf{x}$ as
+Assume that both $\sqrt{k}$ and $\sqrt[4]{k}$ are integers [^4] In fact, we can also construct such examples even though this assumption does not hold.
+However, omitting it will lead to unnecessary complications in the construction (e.g., involving terms like $\lceil\sqrt{k}\rceil$)..
+Construct the target signal $\mathbf{x}$ as
 $|x_{(i)}|^{2}=\begin{cases}\frac{\|\mathbf{x}\|^{2}}{\sqrt[4]{k^{3}}},&i=1,2,\cdots,\sqrt[4]{k},\\ \frac{\|\mathbf{x}\|^{2}}{\sqrt{k}-\sqrt[4]{k}}\times\left(\frac{1}{\sqrt[3]{k}}-\frac{1}{\sqrt{k}}\right),&i=\sqrt[4]{k}+1,\cdots,\sqrt{k},\\ \frac{\|\mathbf{x}\|^{2}}{k-\sqrt{k}}\times\left(1-\frac{1}{\sqrt[3]{k}}\right),&i=\sqrt{k}+1,\cdots,k.\end{cases}$
@@
 $=$ $~\Omega\left(ks(\sqrt[4]{k})\log n\right)=\Omega(k^{3/2}\log n).$
-This demonstrate the improvement in Corollary 3.
+This demonstrates the strict improvement offered by Corollary 3.
 V Proof
-The proofs of Theorem 1 and Corollary 13 are mainly based on the analysis for each step of gESP. In the following, we provide several propositions with respect to each step respectively.
+The proofs of Theorem 1 and Corollary 13 rely primarily on the analysis of each step of gESP. In the following, we establish several propositions corresponding to each step, respectively.
   • Step 1: Generate an index set $S^{0}$ corresponding to the largest $p$ diagonal elements of $\mathbf{Z}$.
     Proposition 1.
-    When the number of samples satisfies $m\geq Cp^{2}s^{2}(p)\log n$, it holds with probability exceeding $1-2e^{-cm}$ that
+    Provided the number of samples satisfies $m\geq Cp^{2}s^{2}(p)\log n$, it holds with probability exceeding $1-2n^{-c}$ that
     $\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)},$ (50)
     where $c$ and $C$ are universal constants.
-  • Step 2: Set $\mathbf{e}^{0}$ as the unit maximal eigenvector of $\mathbf{Z}_{S^{0}}$.
+  • Step 2: Set $\mathbf{e}^{0}$ as the unit leading eigenvector of $\mathbf{Z}_{S^{0}}$.
     Proposition 2.
-    Suppose the set $S^{0}$ contains $p^{\prime}$ indices and satisfies (49), then for any constant $0<\mu_{1}<1$, the inequality
-    $|\mathbf{x}^{*}\mathbf{e}^{0}|\geq\mu_{1}\|\mathbf{x}_{S^{0}}\|$ (50)
-    holds with probability at least $1-2e^{-c_{\mu_{1}}m}$ with the proviso that $m\geq{C_{\mu_{1}}}{p^{\prime}s^{2}(p)}\log n$, where $c_{\mu_{1}}$ and $C_{\mu_{1}}$ are constants related to $\mu_{1}$.
+    Suppose the set $S^{0}$ contains $p^{\prime}$ indices and satisfies (50). Then, the inequality
+    $|\mathbf{x}^{*}\mathbf{e}^{0}|\geq\frac{1}{2}\|\mathbf{x}_{S^{0}}\|$ (51)
+    holds with probability at least $1-n^{-c}$ provided that $m\geq{C}{p^{\prime}s^{2}(p)}\log n$, where $c$ and $C$ are constants.
     Remark 5.
     It may seem unusual to introduce a new parameter $p^{\prime}$ in this proposition, given that $S^{0}$ contains $p$ indices in Algorithm 1.
-However, the proposition holds even when $p^{\prime}=p$, making it a generalized version.
+However, the proposition holds even when $p^{\prime}=p$, serving as a generalization.
 We emphasize that this generalization will be useful in our proofs of Corollary 2 and 3, where we do not have access to the optimal choice of $p$.
-  • Step 3: Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{f}=\mathbf{Z}\mathbf{e}^{0}$ in modulus.
+  • Step 3: Generate an index set $S^{1}$ corresponding to the largest $k$ entries of the vector $\mathbf{f}=\mathbf{Z}\mathbf{e}^{0}$ in magnitude.
     Proposition 3.
-    Denote the set $S_{\gamma}$ as
+    Define the set $S_{\gamma}$ as
 […]
@@
 Moreover, suppose that (50) holds and $\mathbf{e}^{0}$ satisfies (51).
-Then, if the number of samples satisfies $m\geq C_{\mu_{1}}\gamma^{-2}{ks(p)}\log n$, it holds with probability at least $1-4e^{-c_{\mu_{1},\gamma}m}$ that
+Then, if the number of samples satisfies $m\geq C\gamma^{-2}{ks(p)}\log n$, it holds with probability at least $1-n^{-c}$ that
     $S^{1}\supseteq S_{\gamma},$
     where $\mu_{1}$ is the hyperparameter given in (51) and $c_{\mu_{1},\gamma}$ is the constant depending on $\mu_{1}$ and $\gamma$.
-In other words, gESP shall select all the indices of $S_{\gamma}$ in Step 3.
-  • Step 4: Set $\mathbf{z}$ as the maximal eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
+In other words, gESP selects all the indices of $S_{\gamma}$ in Step 3.
+Furthermore, we have
+    $\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$ (52)
+    Crucially, if $|x_{(k)}|$ is extremely small (violating (11)), it simply falls outside $S_{\gamma}$.
+This does not invalidate the proof.
+Instead, the “missed” small entries are absorbed into the approximation error term in Proposition 4.
+Since their magnitudes are small, the total error remains within the $\delta$-neighborhood ($\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$), thus removing the need for a hard lower bound on $x_{(k)}$.
+  • Step 4: Set $\mathbf{z}$ as the leading eigenvector of $\mathbf{Z}_{S^{1}}$ with $\|\mathbf{z}\|^{2}=\lambda^{2}\doteq\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
     Proposition 4.
-    Suppose $S^{1}\supseteq S_{\gamma}$. For any constant $0<\delta<1$ satisfying
-    $\delta^{2}\geq 2-2\sqrt{\left(1-\frac{\gamma^{2}}{4}\right)^{3}}+\frac{\gamma^{2}}{4},$ (51)
-    Step 4 produces a signal estimation $\mathbf{z}$ falling into the $\delta$-neighborhood of $\mathbf{x}$, i.e.,
+    Suppose $S^{1}\supseteq S_{\gamma}$. For any constant $0<\delta<1$, Step 4 produces a signal estimate $\mathbf{z}$ falling into the $\delta$-neighborhood of $\mathbf{x}$, i.e.,
     $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$ (53)
-    with probability exceeding $1-2e^{-c_{\delta}m}$ providing that $m\geq{C_{\delta}}{k}\log n$, where $c_{\delta}$ and $C_{\delta}$ are numerical constants depending on $\delta$.
-Now we are proceeding to prove the theorems in Section III.
+    with probability exceeding $1-n^{-c}$ provided that $m\geq C\delta^{-2}k\log n$, where $c$ and $C$ are numerical constants.
+We defer the proofs of Propositions 14 to the Appendix.
+Now we proceed to prove the theorems in Section III.
 Figure 2: An Illustration of the two functions in (57).
 Proof of Theorem 1.
-Collecting all the results in Proposition 14 with $p^{\prime}=p$, we can show that the sample complexity is
+Combining the results in Proposition 14 with $p^{\prime}=p$, we can show that the sample complexity is
 $m=\Omega\left(\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right),$ (54)
-and the overall probability is $1-e^{-cm}$. Since we take the optimal value for $p$ as
+and the overall probability is $1-n^{-c}$. Since we take the optimal value for $p$ as
 $p_{\text{opt}}={\arg\min}_{p\in[k]}\max\left\{p^{2}s^{2}(p),ks(p)\right\},$
@@
 Proof of Corollary 1.
-For each fixed $p\in[k]$, collecting all the results in Proposition 14 with $p^{\prime}=p$, we can show that the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-e^{-cm}$ when
+For each fixed $p\in[k]$, combining the results in Proposition 14 with $p^{\prime}=p$, we can show that the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-n^{-c}$ when
 $m=\Omega\left(\max\left\{p^{2}s^{2}(p),ks(p)\right\}\log n\right).$
 To ensure that at least one estimate $\mathbf{z}$ satisfies $\operatorname*{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$, we can apply the union bound for all $p\in[k]$.
-Then the sample complexity can be reduced to
+Then the sample complexity becomes
 $m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p)\log n,ks(p)\log n\right\}\right),$
-and the probability is at least $1-ke^{-cm}$. It is actually $1-e^{-cm}$ if we take a sufficiently large constant $c$ since $m>k$. Thus the proof is complete. ∎
+and the probability is at least $1-kn^{-c}$. This is at least $1-n^{-c}$ if we take a sufficiently large constant $c$ since $n>k$. Thus the proof is complete. ∎
 (a) Gaussian signal. $k$ = 10.
 (b) 0-11 signal. $k$ = 10.
-(c) Exponential decaying signal. $k$ = 10
-Figure 3: Performance comparison of relative error and fraction of recovered support as a function of sampling ratio.
+(c) Exponential decaying signal. $k$ = 10.
+Figure 3: Performance comparison of relative error and fraction of recovered support as a function of sample ratio.
 Proof of Corollary 2.
@@
 $\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{\|\mathbf{x}\|^{2}}\overset{(a)}{\geq}\frac{\|\mathbf{x}_{S^{\prime}}\|^{2}}{\|\mathbf{x}\|^{2}}\overset{(b)}{\geq}\frac{1}{2s(p)},$ (55)
-where $(a)$ is because $k\geq p$, and $(b)$ holds via applying Proposition 1 on the set $S^{\prime}$ when $\geq Cp^{2}s^{2}(p)\log n$.
+where $(a)$ is due to $k\geq p$, and $(b)$ holds by applying Proposition 1 to the set $S^{\prime}$ when $m$ $\geq Cp^{2}s^{2}(p)\log n$.
 Then, from the results above and in Proposition 24 (with $p^{\prime}=k$), it holds that for any $p\in[k]$, when
 $m=\Omega\left(\max\left\{p^{2}s^{2}(p)\log n,ks^{2}(p)\log n\right\}\right),$ (56)
-the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-e^{-cm}$.
-This result also holds for all $p\in[k]$ since we can take the union bound, and the probability will be $1-ke^{-cm}$.
-It is actually $1-e^{-cm}$ if we take a sufficiently large constant $c$ since $m>k$.
+the final estimate $\mathbf{z}$ falls into the $\delta$-neighborhood of $\mathbf{x}$ with probability exceeding $1-n^{-c}$.
+This result also holds for all $p\in[k]$ since we can take the union bound, and the probability will be $1-kn^{-c}$.
+It is actually $1-n^{-c}$ if we take a sufficiently large constant $c$ since $n>k$.
 Note that the choice of $p\in[k]$ is arbitrary. Therefore, we can choose the best $p$ to make the sample complexity (56) smallest, which becomes
 $m=\Omega\left(\min_{p\in[k]}\max\left\{p^{2}s^{2}(p)\log n,ks^{2}(p)\log n\right\}\right),$ (57)
-The remaining thing is to show that (56) is equivalent to (44).
-As stated following Definition 2, $ps(p)$ is monotonously increasing, and $s(p)$ is monotonously decreasing.
-Both of them are positive.
-Accordingly, we can make an illustrative plot for the two functions in (56); see Fig.
+It remains to show that (57) is equivalent to (44).
+As stated following Definition 2, $ps(p)$ is monotonically increasing, and $s(p)$ is monotonically decreasing.
+Both functions are positive.
+Accordingly, we can illustrate the behavior of the two functions in (57); see Fig.
 2.
@@
 $p=\lceil{\sqrt{k}}\rceil~~~~\text{or}~~~~p=\lceil{\sqrt{k}}\rceil-1.$ (58)
-From Lemma 1, we know that both the cases imply the same sample complexity
+From Lemma 1, we know that both cases imply the same sample complexity
 $m=\Omega\left(ks^{2}(\lceil{\sqrt{k}}\rceil)\log n\right),$ (59)
@@
 Proof of Corollary 3.
-The proof is similar to the proof of Corollary 2.
+The proof is analogous to the proof of Corollary 2.
 Assume any fixed $p\in[\lceil{\sqrt{k}}\rceil]$.
-The sample complexity for Step 1 is still $\Omega(p^{2}s^{2}(p)\log n)$, and applying Proposition 24 on the rest steps with $p^{\prime}=\lceil{\sqrt{k}}\rceil$ finally leads to
+The sample complexity for Step 1 is still $\Omega(p^{2}s^{2}(p)\log n)$, and applying Proposition 24 on the subsequent steps with $p^{\prime}=\lceil{\sqrt{k}}\rceil$ finally leads to
 $m=\Omega\left(\max\left\{p^{2}s^{2}(p),\sqrt{k}s^{2}(p),ks(p)\right\}\log n\right).$ (60)
@@
 VI-A Experimental Setup
-Across all simulations, the signal dimension is fixed at $n=1000$.
+In all simulations, the signal dimension is fixed at $n=1000$.
 The measurement vectors $\{\mathbf{a}_{i}\}_{i=1}^{m}\subset\mathbb{C}^{n}$ are i.i.d.
@@
 The non-zero entries are then generated from one of three models: i) i.i.d.
-standard complex Gaussian, ii) all ones (binary), and iii) exponentially decaying magnitudes.
+standard complex Gaussian, ii) all ones (binary on support), and iii) exponentially decaying magnitudes.
 All reported results are averaged over $1000$ independent trials to mitigate random fluctuations.
 We compare gESP with CoPRAM, TPM, and ESP.
-For a fair comparison, we consider only the initialization stage of CoPRAM, and use the recommended hyperparameters for TPM.
-For gESP, we set the theoretically optimal parameter $p_{opt}$ from (38) for each trail.
+For a fair comparison, we focus on the initialization stage of CoPRAM, and use the recommended hyperparameters for TPM.
+For gESP, we set the theoretically optimal parameter $p_{opt}$ from (38) for each trial.
 The performance metric is the relative error, defined as:
-$\text{Relative Error}=\frac{\|\mathbf{z}-\mathbf{x}\|_{2}}{\|\mathbf{x}\|_{2}},$
+$\text{Relative Error}=\frac{\operatorname*{dist}(\mathbf{z},\mathbf{x})}{\|\mathbf{x}\|_{2}},$
 where $\mathbf{z}$ is the estimated signal and $\mathbf{x}$ is the true signal [3, 26].
 VI-B Performance Comparison
-In this experiment, we vary the sampling ratio $m/n$ from $0.05$ to $1.0$ and record the corresponding relative error.
+In this experiment, we vary the sample ratio $m/n$ from $0.05$ to $1.0$ and record the corresponding relative error.
 As depicted in Figs.
@@
 For this signal class, gESP was configured using the parameter $p$ determined by our theoretically derived optimum.
-Consistent with the predictions of our theory, gESP achieves the best performance, consistently and slightly outperforming ESP and other state-of-the-art methods.
+Consistent with the predictions of our theory, gESP achieves the best performance, consistently outperforming ESP and other state-of-the-art methods, albeit slightly.
 This clear correspondence between our theory and the empirical results underscores the accuracy of our analysis for signals with rich amplitude variations.
 In contrast, the results for binary and exponentially decaying signals, while not perfectly aligned with asymptotic predictions, offer valuable insights into the practical behavior of the algorithms.
-For binary signals, theory suggests the complexity would always be $\Omega(k^{2}\log n)$ no matter what value $p$ is, and thus one might expect similar performance across methods.
+For binary signals, the theory suggests that the complexity remains $\Omega(k^{2}\log n)$ regardless of the value of $p$, leading one to expect similar performance across methods.
 However, we set $p=k$ in gESP, and it demonstrates a surprisingly large performance margin.
-We conjecture that this significant practical advantage stems from the influence of non-asymptotic constant factors that our analysis does not model but which clearly favor the gESP framework.
+We attribute this significant practical advantage to the influence of nonasymptotic constant factors that our analysis does not model but which clearly favor the gESP framework.
 Similarly, for exponentially decaying signals, gESP is marginally outperformed by ESP.
@@
 This does not contradict our theory but rather highlights that while it provides a powerful general guideline, the precise optimal parameter in practice can be influenced by the signal’s specific structure.
+VI-C Robustness to the Choice of Parameter $p$
+A core feature of gESP lies in that it uses the tuning parameter $p$ to balances the captured signal energy against the accumulated perturbation.
+In this subsection, we experimentally investigate the sensitivity of gESP to the choice of $p$ and demonstrate its practical performance.
+We consider a power-law signal with decay rate $\alpha=0.5$ (i.e., $|x_{(i)}|\propto i^{-0.5}$ for $i=1,\ldots,k$).
+This structure of signal is particularly selected as it is a representative “heavy-tailed” case: neither essentially single-spiked nor perfectly flat.
+Such a signal profile effectively challenges the algorithm to balance energy and perturbation, making it an ideal candidate to evaluate the impact of different selection strategies.
+In the following, we test the performance of gESP with different choices of $p$ on this power-law signal:
+  1. 1.
+    The standard single-peak approach $p=1$ (corresponding to ESP).
+  2. 2.
+    The theoretical oracle $p$ obtained by minimizing the bound in Theorem 1.
+  3. 3.
+    The heuristic surrogate $p=\lceil\sqrt{k}\rceil$ as a middle-ground choice.
+  4. 4.
+    The full support selection $p=k$.
+Figure 4: Performance comparison of gESP with different choices of $p$ under the designed power-law signal.
+We set $n=1000,k=100$ with $m/n$ varying from $0.05$ to $1.0$.
+The multi-peak strategies ($p>1$) significantly outperform the single-peak baseline.
+Notably, the algorithm is robust to large $p$, with the full support choice ($p=k$) achieving the best empirical performance.
+The results are shown in Fig.
+4.
+We observe that all multi-peak strategies ($p>1$) significantly outperform the baseline ESP ($p=1$).
+Notably, the empirical performance remains stable or even improves as $p$ increases, with the full support choice ($p=k$) achieving the best performance in this setting.
+This discrepancy between the theoretical oracle and the empirical optimum can be attributed to two main factors.
+First, our theoretical analysis is asymptotic in nature (assuming $m,n,k\to\infty$), whereas the experiments are conducted in a finite-dimensional setting ($n=1000,k=100$).
+Clearly the finite-sample behavior may arguably tolerate more perturbation than the asymptotic worst-case prediction.
+Second, the “oracle” choice is derived by minimizing a theoretical upper bound on the sample complexity.
+This bound relies on standard concentration inequalities (e.g., Bernstein-type bounds), which are sufficient yet inherently conservative.
+As a result, the theoretical derivation tends to overestimate the perturbation accumulation associated with large $p$, leading to a more conservative recommendation compared to the algorithm’s actual robustness behavior in practice.
+Practically, this result is encouraging: it implies that users can simply select a sufficiently large $p$ (e.g., above the expected sparsity level) to achieve the near-optimal performance without precise knowledge of the energy profile $s(p)$.
 VII Conclusion
-In this paper, we investigate the problem of sparse phase retrieval, focusing on improving initialization algorithms for recovering a $k$-sparse $n$-dimensional signal from $m$ phaseless observations.
-Through an analysis of existing methods, we propose generalized Exponential Spectral Pursuit (gESP), which enhances initialization performance and refines the understanding of sample complexity in sparse phase retrieval.
+In this paper, we have addressed the problem of sparse phase retrieval, focusing on improving initialization algorithms for recovering a $k$-sparse $n$-dimensional signal from $m$ phaseless observations.
+Motivated by the limitations of existing methods that rely on specific structural assumptions of signals, we have proposed generalized Exponential Spectral Pursuit, which promotes the initialization performance and also enhances the understanding of sample complexity in sparse phase retrieval.
 Empirical evaluations further validate the practical effectiveness of gESP.
-Simulation experiments demonstrated its robustness and efficiency in recovering sparse signals, highlighting its potential utility in real-world applications.
-While this work advances both theoretical bounds and practical performance for sparse phase retrieval, it does not fully resolve the statistical-to-computational gap.
-Specifically, under Gaussian measurements, $\Omega(k\log n)$ samples are theoretically sufficient to determine the target signal.
-However, in practical scenarios requiring polynomial-time algorithms, the sample complexity increases to $\Omega(k^{2}\log n)$ in the worst case, such as when the target signal consists of 0-11 entries.
+Simulation experiments demonstrate its robustness and efficiency in recovering sparse signals, highlighting its potential utility in real-world applications.
+Although this work advances both theoretical bounds and practical performance for sparse phase retrieval, it does not fully resolve the statistical-to-computational gap.
+Specifically, under Gaussian measurements, an information-theoretic bound of $\Omega(k\log n)$ samples is sufficient to uniquely determine the target signal.
+However, for polynomial-time recovery, the sample complexity increases to $\Omega(k^{2}\log n)$ in the worst case, for example, when the target signal consists of binary entries.
 Future research could aim to bridge this gap by exploring novel algorithmic frameworks or leveraging insights from other areas, such as computational complexity and optimization.
-Additionally, extending the applicability of gESP to broader measurement models and refining the refinement stage offer promising directions for advancing this field.
+Furthermore, extending the applicability of gESP to broader measurement models and optimizing the subsequent refinement stage offer promising directions to advance this field.
 Appendix A Basic Tools and Lemmas
-In this section, we present several useful lemmas and necessary proofs. First of all, recall that
+In this section, we present several useful lemmas and necessary proofs. Recall the definition of $\mathbf{Z}$:
 $\mathbf{Z}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right)\mathbf{a}_{i}\mathbf{a}_{i}^{*},$ (A.1)
-where $\lambda^{2}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$. It is challenging to derive the expectation of $\mathbf{Z}$. Following the technique in [12, 28], we introduce
+where $\lambda^{2}=\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}$.
+Deriving the exact expectation of $\mathbf{Z}$ is challenging due to the dependency between $y_{i}$ and $\lambda^{2}$.
+Following the technique in [31, 33], we introduce a proxy matrix:
 $\tilde{\mathbf{Z}}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{a}_{i}\mathbf{a}_{i}^{*},$ (A.2)
@@
 $\mathbf{f}$ $=\mathbf{Z}\mathbf{e}^{0},$ $\tilde{\mathbf{f}}$ $=\tilde{\mathbf{Z}}\mathbf{e}^{0}.$
-In the following, we introduce several results concentrating on the effect of the ceiling sign on the sample complexity.
+In the following, we introduce several results concerning the effect of the ceiling function on the sample complexity.
 Lemma 1.
@@
 $\Omega(\lceil\sqrt{k}\rceil)$ $=\Omega(\sqrt{k}),$ (A.5)
-$\Omega(s(\lceil\sqrt{k}\rceil)$ $=\Omega(s(c\lceil\sqrt{k}\rceil)),$ (A.6)
+$\Omega(s(\lceil\sqrt{k}\rceil))$ $=\Omega(s(c\lceil\sqrt{k}\rceil)),$ (A.6)
 where $c$ is a fixed positive integer.
 Proof.
-Note that $\sqrt{k}\leq\lceil\sqrt{k}\rceil\leq\sqrt{k}+1$, then (A.5) can be easily derived from the following relationship.
+Note that $\sqrt{k}\leq\lceil\sqrt{k}\rceil\leq\sqrt{k}+1$, then (A.5) can be easily derived from the limit:
 $\lim_{k\to\infty}\frac{\sqrt{k}+1}{\sqrt{k}}=1.$
@@
 $s(c\lceil\sqrt{k}\rceil)=\frac{\|\mathbf{x}\|^{2}}{\sum_{j=1}^{c\lceil\sqrt{k}\rceil}|x_{(j)}|^{2}}.$
-Since $x_{(j)}$’s are the rearrangement of the entries of $\mathbf{x}$ with a descend order in their modulus, the summation in the denominator can be estimated as
+Since $x_{(j)}$’s are the rearrangement of the entries of $\mathbf{x}$ in descending order of their magnitudes, the summation in the denominator can be estimated as
 $\sum_{j=1}^{\lceil\sqrt{k}\rceil}|x_{(j)}|^{2}\leq\sum_{j=1}^{c\lceil\sqrt{k}\rceil}|x_{(j)}|^{2}\leq c\sum_{j=1}^{\lceil\sqrt{k}\rceil}|x_{(j)}|^{2},$
@@
 $\frac{1}{c}s(\lceil\sqrt{k}\rceil)\leq s(c\lceil\sqrt{k}\rceil)\leq s(\lceil\sqrt{k}\rceil).$
-In that $c$ is a constant as $k$ and $n$ tend to infinity, (A.6) is a natural derivation. ∎
-Lemma 2 (Bernstein inequality for complex cases).
-Let $X_{i}$ be independent, mean-zero complex sub-exponential random variables for $i\in[m]$, then for any $\epsilon>0$, we have
-$\mathbb{P}\left(\left\lvert\frac{1}{m}\sum_{i=1}^{m}X_{i}\right\rvert\geq\epsilon\right)\leq 4\exp\left(-cm\min\left\{\frac{\epsilon^{2}}{K^{2}},\frac{\epsilon}{K}\right\}\right),$
-where $K=\mathop{\max}_{i\in[m]}\|X_{i}\|_{\psi_{1}}$ and $c$ is the numerical constant. Also see [22, Lemma 10].
-Next, the difference between $\lambda^{2}$ and $\|\mathbf{x}\|^{2}$ is analyzed.
+Since $c$ is a constant as $k$ and $n$ tend to infinity, (A.6) follows naturally. ∎
+The next lemma gives the error bound of $\lambda^{2}$.
 Lemma 2.
-Given $k$-sparse signal $\mathbf{x}$, for any constant $\beta>0$, suppose that $\mathbf{a}_{i}$’s are complex Gaussian random variables. Then provided that $m\geq C_{\beta}\kappa^{2}$,
-$\left(1-\frac{\beta}{\kappa}\right)\|\mathbf{x}\|^{2}\leq\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}\leq\left(1+\frac{\beta}{\kappa}\right)\|\mathbf{x}\|^{2}$
-holds with probability at least $1-e^{-c_{\beta}m}$, where $\kappa$ can be substitute by any quantity satisfying $\frac{\beta}{\kappa}<1$, such as $ps(p)$, and $c_{\beta}$ and $C_{\beta}$ are numerical constants determined by $\beta$.
+Consider the problem (1). With probability at least $1-n^{-c}$, it holds that
+$\left|\lambda^{2}-\|\mathbf{x}\|^{2}\right|\leq C\sqrt{\frac{\log n}{m}}\|\mathbf{x}\|^{2}.$
 Proof.
-It is easy to find that $|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}$’s are independent sub-exponential random variables with mean $\mathbb{E}[|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}]=\|\mathbf{x}\|^{2}$ and $c_{0}\|\mathbf{x}\|^{2}$ as their squared sub-exponential norms.
-Hence, employing Bernstein’s inequality yields
-$\mathbb{P}\left(\left\lvert\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}-\|\mathbf{x}\|^{2}\right\rvert>\frac{\beta}{\kappa}\|\mathbf{x}\|^{2}\right)$
-$~~~~~~~\leq 2\exp\left(-C_{\beta}m\min\left\{\frac{\beta^{2}}{\kappa^{2}K^{2}},\frac{\beta}{\kappa K}\right\}\right),$
-where $K=\mathop{\max}_{i\in[m]}\||\mathbf{a}_{i}^{*}\mathbf{x}|^{2}\|_{\psi_{1}}/\|\mathbf{x}\|^{2}$ and $C_{\beta}$ is the numerical constant. Thus we obtain the desired result.∎
-Finally, we introduce a useful lemma that measures the similarity between the exponential spectral $\mathbf{Z}$ and its estimator $\mathbb{E}[\tilde{\mathbf{Z}}$].
+Let $y_{i}=\mathbf{a}_{i}^{*}\mathbf{x}$.
+Since $\mathbf{a}_{i}\sim\mathcal{CN}(0,\mathbf{I})$, the variable $y_{i}\sim\mathcal{CN}(0,\|\mathbf{x}\|^{2})$.
+The term $|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}/\|\mathbf{x}\|^{2}$ follows a standard exponential distribution, which is sub-exponential with constant norm.
+We apply the standard Bernstein inequality for the sum of independent sub-exponential random variables [23, Theorem 2.8.1]. For any $t\geq 0$,
+$\mathbb{P}\left(\left|\frac{1}{m}\sum_{i=1}^{m}|\mathbf{a}_{i}^{*}\mathbf{x}|^{2}-\|\mathbf{x}\|^{2}\right|\geq t\|\mathbf{x}\|^{2}\right)$
+$~~~~~~~~~~\leq 2\exp\left(-cm\min(t^{2},t)\right).$
+Setting $t=C\sqrt{\frac{\log n}{m}}$ yields the desired result. ∎
+The following lemma gives the operator norm bound of the weighted average of random matrices, which is a natural extension of the classical result in random matrix theory.
+Lemma 3 (Uniform Concentration with Bounded Weights).
+Let $\mathbf{a}_{i}\sim\mathcal{CN}(0,\mathbf{I}_{n})$ be independent standard complex Gaussian vectors.
+Let $\xi_{i}\in\mathbb{R}$ be independent scalar random variables (potentially depending on $\mathbf{a}_{i}$) satisfying $|\xi_{i}|\leq B$ almost surely for some constant $B>0$.
+With probability at least $1-2\exp(-cs\log(n/s))$, the following holds uniformly for all subsets $S\subset[n]$ with cardinality $|S|=s$, provided the sample size satisfies $m\geq Cs\log(n/s)$:
+$\left\|\frac{1}{m}\sum_{i=1}^{m}\xi_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}-\mathbb{E}\left[\xi_{1}(\mathbf{a}_{1})_{S}(\mathbf{a}_{1})_{S}^{*}\right]\right\|$
+$\leq$~ $C^{\prime}B\sqrt{\frac{s\log(n/s)}{m}},$ (A.7)
+where $c,C,C^{\prime}$ are universal numerical constants.
+Proof.
+The proof follows the standard covering argument used for covariance matrix estimation (see, e.g., the proof of [23, Theorem 4.6.1]).
+The basic intuition is that the weighted bounded scalar will not essentially change the sub-exponential property of the random matrix.
+For simplicity, we denote
+$\mathbf{E}_{i}:=\xi_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}-\mathbb{E}\left[\xi_{1}(\mathbf{a}_{1})_{S}(\mathbf{a}_{1})_{S}^{*}\right],$
+and $\mathbf{E}:=\frac{1}{m}\sum_{i=1}^{m}\mathbf{E}_{i}$.
+Step 1: Discretization via $\epsilon$-net.
+Fix a support $S$ with $|S|=s$.
+Let $\mathcal{N}$ be a $\frac{1}{4}$-net of the unit sphere in $\mathbb{C}^{|S|}$.
+The cardinality of the net is bounded by $|\mathcal{N}|\leq 9^{2|S|}\leq 9^{4s}$.
+From [23, Lemma 4.4.1], the operator norm of a Hermitian matrix $\mathbf{E}$ is controlled by the net
+$\|\mathbf{E}\|\leq 2\sup_{u\in\mathcal{N}}|\mathbf{u}^{*}\mathbf{E}\mathbf{u}|.$ (A.8)
+Thus, it suffices to bound the scalar deviation for each fixed $\mathbf{u}\in\mathcal{N}$.
+Step 2: Scalar Concentration. Fix a unit vector $\mathbf{u}\in\mathcal{N}$. Consider the scalar random variable
+$Z_{i}=\mathbf{u}^{*}\left(\xi_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right)\mathbf{u}=\xi_{i}|\langle(\mathbf{a}_{i})_{S},\mathbf{u}\rangle|^{2}.$
+Since $(\mathbf{a}_{i})_{S}$ is standard Gaussian, the inner product $g=\langle(\mathbf{a}_{i})_{S},\mathbf{u}\rangle$ is a standard scalar complex Gaussian variable.
+The term $|g|^{2}$ follows an exponential distribution, which is sub-exponential with norm $\||g|^{2}\|_{\psi_{1}}\leq C_{sub}$.
+Because the weights are bounded ($|\xi_{i}|\leq B$), the variable $Z_{i}$ is sub-exponential with norm
+$\|Z_{i}\|_{\psi_{1}}=\|\xi_{i}|g|^{2}\|_{\psi_{1}}\leq B\||g|^{2}\|_{\psi_{1}}\leq C_{sub}B.$
+We now apply the scalar Bernstein inequality for sub-exponential variables [23, Theorem 2.8.1] and obtain
+$\mathbb{P}\left(\left|\frac{1}{m}\sum_{i=1}^{m}(Z_{i}-\mathbb{E}Z_{i})\right|\geq t\right)$
+$~~~~~~~~\leq 2\exp\left(-cm\min\left(\frac{t^{2}}{K^{2}},\frac{t}{K}\right)\right),$
+where $K=\max_{i}\|Z_{i}\|_{\psi_{1}}\lesssim B$. Assuming large $m$ such that the sub-gaussian tail dominates, we set $t=C^{\prime}B\sqrt{\frac{s\log(n/s)}{m}}$ and get
+$\mathbb{P}\Big(\Big|\underbrace{\frac{1}{m}\sum_{i=1}^{m}(Z_{i}-\mathbb{E}Z_{i})}_{\mathbf{u}^{*}\mathbf{E}_{S}\mathbf{u}}\Big|\geq t\Big)\leq~2\exp(-c^{\prime}s\log(n/s)).$
+Note that the $B^{2}$ in the variance term cancels out, leaving the rate dependent on $m$ and $s$.
+Step 3: Union Bound.
+We now control the probability that the operator norm exceeds $t$ for any valid support $S$.
+Taking the union bound over all $\binom{n}{s}$ supports and all $\mathbf{u}\in\mathcal{N}$ for each support yields
+[…]
+$\makebox[61.83356pt][c]{$\leq$}\binom{n}{s}\cdot 9^{4s}\cdot 2\exp\left(-c^{\prime}s\log(n/s)\right)$
+$\makebox[61.83356pt][c]{$\leq$}\left(\frac{en}{s}\right)^{s}9^{4s}\cdot 2\exp\left(-c^{\prime}s\log(n/s)\right).$
+By choosing the constant $C$ in the sample complexity $m\geq Cs\log(n/s)$ sufficiently large, the exponent $c^{\prime}s\log(n/s)$ dominates and thus ensures that the total probability is at most $2\exp(-cs\log(n/s))$.
+∎
+Based on the above lemmas, we proceed to show the similarity between the exponential spectral $\mathbf{Z}$ and its estimator $\mathbb{E}[\tilde{\mathbf{Z}}$].
+First, we bound the operator norm of the difference matrix $\tilde{\mathbf{Z}}-\mathbb{E}[\tilde{\mathbf{Z}}$].
+In fact, this is a direct application of Lemma 3.
 Lemma 4.
-Suppose a constant $\eta>0$ and a positive integer $q\leq k$.
-For all $S\subseteq[n]$ satisfying $|S|=q$, let $T$ be a subset of $\mathbb{C}^{n}$ such that $\|\mathbf{z}\|\leq 1$ holds for all $\mathbf{z}\in T$.
-Given any constant $\epsilon\in(0,1)$, let $\mathcal{N}(T,\epsilon)$ be the covering number [31, Definition 4.2.2] of $T$ and $g(k,s(p))$ be a function with respect to $k$ and $s(p)$ satisfying $g(k,s(p))\geq 1$.
-It holds with probability at least $1-2e^{-c_{\eta}m}$ that
-$\sup_{\mathbf{z}_{1},\mathbf{z}_{2}\in T}\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{2}\leq\frac{\eta}{g(k,s(p))}$ (A.7)
-with the condition that
-$m\geq C_{\eta}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)),$ (A.8)
-where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
-In particular, when $T=\{\mathbf{u}|\operatorname{supp}(\mathbf{u})=S,\|\mathbf{u}\|\leq 1\}$, the inequality
-$\left\lVert\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right\rVert\leq\frac{\eta}{2g(k,s(p))}$ (A.9)
-holds true given $m\geq C_{\eta}qg^{2}(k,s(p))\log n$.
+Let $\tilde{\mathbf{Z}}$ be defined as in (A.2) and its expectation is given by (A.3).
+For all $S\subseteq[n]$ with $|S|=s$, the following bound holds for the operator norm with probability exceeding $1-2\exp(-cs\log(n/s))$:
+$\|\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\|\leq C\sqrt{\frac{s\log(n/s)}{m}}.$ (A.9)
 Proof.
-Denote the smallest $\epsilon$-net of $T$ as $N_{\epsilon}^{T}$, which means that $|N_{\epsilon}^{T}|=\mathcal{N}(T,\epsilon)$.
-Then, for any unit $\mathbf{z}_{1},\mathbf{z}_{2}\in T$, there exist $\mathbf{z}_{10}\in N_{\epsilon}^{T}$ and $\mathbf{z}_{20}\in N_{\epsilon}^{T}$ such that $\|\mathbf{z}_{1}-\mathbf{z}_{10}\|\leq\epsilon$ and $\|\mathbf{z}_{2}-\mathbf{z}_{20}\|\leq\epsilon$.
-It holds that
-$\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{2}$
-== $\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}+\mathbf{z}_{1}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)(\mathbf{z}_{2}-\mathbf{z}_{20})$
-$+(\mathbf{z}_{1}-\mathbf{z}_{10})^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}$
-$\overset{(a)}{\leq}$ $\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right)\mathbf{z}_{20}+2\epsilon\left\lVert\mathbf{Z}_{S}-{\mathbb{E}[\tilde{\mathbf{Z}}_{S}]}\right\rVert,$
-where $(a)$ comes from the definition of $N_{\epsilon}^{T}$ and the matrix norm, together with the fact that $\|\mathbf{z}_{20}\|\leq 1$. We first prove the following two inequalities
-$\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))},$ (A.10)
-$\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))}.$ (A.11)
-$\dagger$} Proof for (A.10)
-Simple calculation gives that
-$\mathbf{z}_{10}^{*}\tilde{\mathbf{Z}}_{S}\mathbf{z}_{20}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{\mathbf{y}_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20},$
-which is the sum of independent sub-exponential random variables with constant $\psi_{1}$ norms, since $\|\mathbf{z}_{10}\|,\|\mathbf{z}_{20}\|\leq 1$. Employing Bernstein’s inequality yields
-$\mathbb{P}\left(\left\lvert\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\right\rvert>\frac{\eta}{4}\times\frac{1}{g(k,s(p))}\right)$
-$\leq$~ $2\exp\left(-C^{\prime}_{\eta}m\min\left\{\frac{\eta^{2}}{16K^{2}g^{2}(k,s(p))},\frac{\eta}{4Kg(k,s(p))}\right\}\right)$
-$\doteq$~ $p_{1},$
-where the sub-exponential norm satisfies
-$K=\max_{i\in[m]}\left\rVert\left(\frac{1}{2}-\exp\left(-\frac{\mathbf{y}_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\left\lvert(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{10}\right\rvert\left\lvert(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert\right\rVert_{\psi_{1}}$
-and $c_{\eta}$ is the numerical constant.
-Take the union bound for all $\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}$, we have
-$~\mathbb{P}\left(\sup_{\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}}\mathbf{z}_{10}^{*}\left(\tilde{\mathbf{Z}}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\right)\mathbf{z}_{20}\leq\frac{\eta}{4}\times\frac{1}{g(k,s(p))}\right)$
-$\geq$ $~1-\mathcal{N}^{2}(T,\epsilon)\times p_{1}.$
-Furthermore, take the union bound for all $S\subseteq[n]$ satisfying $|S|\leq q$, the probability will exceed $1-{n\choose q}\mathcal{N}^{2}(T,\epsilon)p_{1}$.
-Therefore, when
-$m\geq C_{\eta}^{\prime}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)),$
-(A.10) holds with probability exceeding $1-2\exp\big(c^{\prime}_{\eta}m\big)$.
-$\dagger$} Proof for (A.11)
-Direct calculation yields
-$\left\rvert\mathbf{z}_{10}^{*}\left(\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\right)\mathbf{z}_{20}\right\rvert$
-== $\left\lvert\frac{1}{m}\sum_{i=1}^{m}\left(\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)\right)\mathbf{z}_{10}^{*}(\mathbf{a}_{i}\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$ (A.14)
-$\overset{(a)}{=}$ $\max_{i\in[m]}\exp\left(-\frac{y_{i}^{2}}{\xi}\right)\hskip-0.7113pt\frac{y_{i}^{2}}{\xi^{2}}\hskip-0.7113pt\big|\lambda^{2}-\|\mathbf{x}\|^{2}\big|\hskip-0.7113pt\hskip-0.7113pt\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i}\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$ (A.17)
-$\overset{(b)}{\leq}$ $~\frac{1}{\xi}\big|\lambda^{2}-\|\mathbf{x}\|^{2}\big|\times\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert$, (A.20)
-where $(a)$ comes from Lagrange’s mean value formula with $F(t)=\exp\big(\frac{y_{i}^{2}}{t}\big)$ and $\xi\in\left[\min\{\lambda^{2},\|\mathbf{x}\|^{2}\},\max\{\lambda^{2},\|\mathbf{x}\|^{2}\}\right]$, while $(b)$ employs the numerical inequality $xe^{-x}\leq 1,\ x\geq 0$.
-For the second term, since $\mathbf{a}_{i}$’s are complex standard Gaussian random vectors and $\|\mathbf{z}_{10}\|,\|\mathbf{z}_{20}\|\leq 1$, it can be derived that for every $i\in[m]$, $\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}$ is sub-exponential random variable with constant $\psi_{1}$ norm and expectation being $(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}$.
-Therefore, follow the same concentration analysis as in the proof of (A.10), we conclude that the inequality
-$\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}-(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}\right\rvert\leq\frac{\eta}{12}\times\frac{1}{g(k,s(p))}$ (A.21)
-holds for all $\mathbf{z}_{10},\mathbf{z}_{20}\in N_{\epsilon}^{T}$ and $S\subseteq[n]$ satisfying $|S|\leq q$ with probability exceeding $1-2\exp(c^{\prime\prime}_{\eta}m)$ when
-$m\geq C^{\prime\prime}_{\eta}g^{2}(k,s(p))(\log\mathcal{N}(T,\epsilon)+q\log(n/q)).$
-Note that
-$\left\lvert(\mathbf{z}_{10}^{*})_{S}(\mathbf{z}_{20})_{S}\right\rvert\leq\sqrt{\|\mathbf{z}_{10}\|^{2}\|\mathbf{z}_{20}\|^{2}}\leq 1.$
-The inequality (A.21) can be transformed to
-$\left\lvert\frac{1}{m}\sum_{i=1}^{m}\mathbf{z}_{10}^{*}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i}^{*})_{S}\mathbf{z}_{20}\right\rvert\leq 1+\frac{\eta}{12}\times\frac{1}{g(k,s(p))}.$ (A.22)
-For the first term, by employing Lemma 3 with with $\kappa=g(k,s(p))$ and $\beta=\eta/12$, we conclude that
-$\frac{1}{\xi}\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert$ $\leq\frac{1}{\|\mathbf{x}\|^{2}-\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert}\left\lvert\lambda^{2}-\|\mathbf{x}\|^{2}\right\rvert$
-$\leq\frac{1}{\left(1-\frac{\eta}{12g(k,s(p))}\right)\|\mathbf{x}\|^{2}}\times\frac{\eta\|\mathbf{x}\|^{2}}{12g(k,s(p))}$ (A.23)
-holds true with probability exceeding $1-e^{-cm}$ as soon as
-$m\geq C^{\prime\prime}_{\eta}g^{2}(k,s(p)).$
-Finally, taking (A.22) and (A.23) into (A.20) yields
-[…]
-$\overset{(a)}{\leq}$ $~3\times\frac{\eta}{12}\times\frac{1}{g(k,s(p))}=\frac{\eta}{4}\times\frac{1}{g(k,s(p))},$
-where $(a)$ comes from the property $\frac{1+x}{1-x}\leq 3$ with $x\leq\frac{1}{2}$ and the fact that $\frac{\eta}{12}\times\frac{1}{g(k,s(p))}\leq\frac{1}{2}$.
-Here, we complete the proof of two sub-conclusions.
-Finally, combining (A.10) and (A.11) with $T=\{\mathbf{u}|\operatorname{supp}(\mathbf{u})=S,\|\mathbf{u}\|\leq 1\}$ yields (A.9), which in turn proves our conclusion.
-∎
-Remark 4.
-As indicated above, the term $q\log(n/q)$ arises from the union bound for $S\in[n],|S|=q$. For fixed set $S$, the sample complexity reduces to
-$m\geq C_{\eta}g^{2}(k,s(p))\log\mathcal{N}(T,\epsilon).$ (A.24)
+Note that the scalar coefficients $\xi_{i}=\frac{1}{2}-\exp(-y_{i}^{2}/\|\mathbf{x}\|^{2})$ are strictly bounded, since $|\xi_{i}|\leq 1/2$. By applying Lemma 3, we obtain the desired result. ∎
+Next, we bound the operator norm of the difference matrix $\mathbf{Z}-\tilde{\mathbf{Z}}$.
+Lemma 5.
+Let $\mathbf{Z}$ and $\tilde{\mathbf{Z}}$ be defined as in (A.1) and (A.2).
+For all $S\subseteq[n]$ with $|S|=s$, the following bound holds for the operator norm with probability exceeding $1-2\exp(-cs\log(n/s))$:
+$\|\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}\|\leq C\sqrt{\frac{\log n}{m}}.$ (A.10)
+Proof.
+Write the difference matrix $\Delta\mathbf{Z}_{S}=\mathbf{Z}_{S}-\tilde{\mathbf{Z}}_{S}$:
+$\Delta\mathbf{Z}_{S}=\frac{1}{m}\sum_{i=1}^{m}\left[\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right](\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}.$ (A.11)
+Define the scalar function $g(t)=\exp(-y_{i}^{2}/t)$.
+By the Lagrange Mean Value Theorem, there exists a $\xi_{i}$ between $\min(\lambda^{2},\|\mathbf{x}\|^{2})$ and $\max(\lambda^{2},\|\mathbf{x}\|^{2})$ such that
+$g(\|\mathbf{x}\|^{2})-g(\lambda^{2})=g^{\prime}(\xi_{i})(\|\mathbf{x}\|^{2}-\lambda^{2}).$
+The derivative is $g^{\prime}(t)=\frac{y_{i}^{2}}{t^{2}}\exp(-y_{i}^{2}/t)$. Thus,
+$\exp\left(-\frac{y_{i}^{2}}{\|\mathbf{x}\|^{2}}\right)-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)=\underbrace{\frac{y_{i}^{2}}{\xi_{i}^{2}}\exp\left(-\frac{y_{i}^{2}}{\xi_{i}}\right)}_{\delta_{i}\geq 0}(\|\mathbf{x}\|^{2}-\lambda^{2}).$
+Substituting this back into the matrix sum yields
+$\Delta\mathbf{Z}_{S}=(\|\mathbf{x}\|^{2}-\lambda^{2})\cdot\left(\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right).$
+Now, we apply the operator norm $\|\cdot\|$. Since $|\lambda^{2}-\|\mathbf{x}\|^{2}|$ is a global scalar, we can pull it out and obtain
+$\|\Delta\mathbf{Z}_{S}\|=|\|\mathbf{x}\|^{2}-\lambda^{2}|\cdot\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|.$ (A.12)
+First, we bound the scalar coefficient $\delta_{i}$ uniformly. Consider the function $h(u)=ue^{-u}$ for $u\geq 0$. Its maximum value is $e^{-1}$ at $u=1$.We can rewrite $\delta_{i}$ as
+$\delta_{i}=\frac{1}{\xi_{i}}\left(\frac{y_{i}^{2}}{\xi_{i}}\exp\left(-\frac{y_{i}^{2}}{\xi_{i}}\right)\right)\leq\frac{1}{\xi_{i}}\cdot e^{-1}.$
+Note that for each $i$, $\xi_{i}$ is between $\min(\lambda^{2},\|\mathbf{x}\|^{2})$ and $\max(\lambda^{2},\|\mathbf{x}\|^{2})$.
+From Lemma 2, denote $\epsilon=C\sqrt{\frac{\log n}{m}}$, and we have $\xi_{i}\geq(1-\epsilon)\|\mathbf{x}\|^{2}$.
+Thus,
+$0\leq\delta_{i}\leq\frac{1}{e(1-\epsilon)\|\mathbf{x}\|^{2}}\doteq\frac{C_{1}}{\|\mathbf{x}\|^{2}}.$ (A.13)
+Second, we bound $\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|$ via the PSD property.
+Since $(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}$ are Positive Semidefinite (PSD) matrices and coefficients $\delta_{i}$ are non-negative, it holds that
+$\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|\leq\left(\max_{i}\delta_{i}\right)\left\|\frac{1}{m}\sum_{i=1}^{m}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|.$
+Using the bound (A.13) for $\delta_{i}$ and applying Lemma 3 with $\xi_{i}\equiv 1$, we have
+$\left\|\frac{1}{m}\sum_{i=1}^{m}\delta_{i}(\mathbf{a}_{i})_{S}(\mathbf{a}_{i})_{S}^{*}\right\|\leq\frac{C_{1}}{\|\mathbf{x}\|^{2}}\left(1+C_{2}\sqrt{\frac{s\log(n/s)}{m}}\right).$
+Finally, substituting this back into equation (A.12) yields
+$\|\Delta\mathbf{Z}_{S}\|\leq|\|\mathbf{x}\|^{2}-\lambda^{2}|\cdot\frac{C_{1}}{\|\mathbf{x}\|^{2}}\left(1+C_{2}\sqrt{\frac{s\log(n/s)}{m}}\right).$
+Applying Lemma 2, we have
+$\|\Delta\mathbf{Z}_{S}\|\leq C_{1}C_{3}\sqrt{\frac{\log n}{m}}\cdot\left(1+C_{2}\sqrt{\frac{s\log(n/s)}{m}}\right).$
+Choosing $m\geq C_{4}s\log(n/s)$ with sufficiently large $C_{4}$ allows the union bound to control the failure probability by $2\exp(-cs\log(n/s))$. This completes the proof. ∎
+As a consequence, via triangle inequality, we have the following result.
+Corollary 4.
+For all $S\subseteq[n]$ with $|S|=s$, the following concentration inequality
+$\|\mathbf{Z}_{S}-\mathbb{E}[\tilde{\mathbf{Z}}_{S}]\|\leq C\sqrt{\frac{s\log(n/s)}{m}}$ (A.14)
+holds true with probability exceeding $1-2e^{-cs\log(n/s)}$ given $m\geq Cs\log(n/s)$, where $c$ and $C$ are numerical constants.
 A-A Concentration analysis for the sum of the diagonal entries
 Lemma 6.
-For all $\mathcal{S}\subseteq[n]$ with $|\mathcal{S}|=p$ and constant $0<\eta<1$, the following concentration inequality
-$\left\lvert\frac{1}{p}\sum_{j\in\mathcal{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert<\eta\times\frac{1}{ps(p)}$ (A.25)
-holds true with probability exceeding $1-2e^{-c_{\eta}m}$ given $m\geq C_{\eta}p^{2}s^{2}(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
+For all $S\subseteq[n]$ with $|S|=p$ and constant $0<\eta<1$, the following concentration inequality
+$\left\lvert\frac{1}{p}\sum_{j\in S}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert<\eta\times\frac{1}{ps(p)}$ (A.15)
+holds true with probability exceeding $1-n^{-c_{\eta}}$ given $m\geq C_{\eta}p^{2}s^{2}(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constants related to $\eta$.
 Proof.
-We can see from (A.3) that
-$\mathbb{E}[\tilde{Z}_{j}]=\left\{\begin{array}[]{cc}0,&j\notin{\text{supp}(\mathbf{x})},\\ \frac{|x_{j}|^{2}}{4\|\mathbf{x}\|^{2}},&j\in{\text{supp}(\mathbf{x})}.\end{array}\right$. (A.26)
-Therefore, the left term of (A.25) can be expressed as
-$\frac{1}{p}\sum_{j\in\mathcal{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)=\frac{1}{p}\sum_{j\in\mathcal{S}}\left(\mathbf{Z}_{\{j\}}\hskip-0.7113pt-\hskip-0.7113pt\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right).$
-For all $j\in[n]$, employing Lemma 4 with $S=\{j\}$, the fixed $T=\{\mathbf{1},-\mathbf{1}\}$, and $g(k,s(p))=ps(p)$, we can derive that
-$\sup_{j\in[n]}\mathbf{1}^{*}\left(\mathbf{Z}_{\{j\}}-\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right)\mathbf{1}\leq\eta\times\frac{1}{ps(p)},$
-$\sup_{j\in[n]}\mathbf{1}^{*}\left(\mathbf{Z}_{\{j\}}-\frac{\mathbf{x}_{\{j\}}\mathbf{x}_{\{j\}}^{*}}{4\|\mathbf{x}\|^{2}}\right)(-\mathbf{1})\leq\eta\times\frac{1}{ps(p)}.$
-holds with probability exceeding $1-2e^{-c_{\eta}m}$ and the corresponding sample complexity becomes
-$m\geq C_{\eta}p^{2}s^{2}(p)\log n,$
-which in turn proves (A.25) since $\mathcal{S}\subseteq[n]$.
+The left term of (A.15) can be expressed as
+$\frac{1}{p}\sum_{j\in S}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)$ $=\frac{1}{p}\sum_{j\in S}\left(\mathbf{Z}_{\{j\}}-\mathbb{E}[\tilde{\mathbf{Z}}_{\{j\}}]\right)$
+$\leq\max_{j\in S}\left\lvert\mathbf{Z}_{\{j\}}-\mathbb{E}[\tilde{\mathbf{Z}}_{\{j\}}]\right\rvert$.
+For all $j\in[n]$, employing Corollary 4 with $S=\{j\}$ (and thus $|S|=1$), we have
+$\max_{j\in S}\left\lvert\mathbf{Z}_{\{j\}}-\mathbb{E}[\tilde{\mathbf{Z}}_{\{j\}}]\right\rvert\leq C\sqrt{\frac{\log n}{m}}.$
+Taking $m\geq C\eta^{-2}p^{2}s^{2}(p)\log n$ yields the desired result.
 A-B Concentration analysis for $\mathbf{f}$ in (A.4)
@@
 $\max_{l\in[n]}\left\lvert f_{l}-\mathbb{E}(\tilde{f}_{l})\right\rvert<\eta\times\frac{\gamma}{2\sqrt{ks(p)}}$ (A.16)
-holds with probability exceeding $1-2\exp(-c_{\eta}m)$ given $m\geq C_{\eta}\frac{ks(p)}{\gamma^{2}}\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constant related to $\eta$.
+holds with probability exceeding $1-n^{-c_{\eta}}$ given $m\geq C_{\eta}{\gamma^{-2}}ks(p)\log n$, where $c_{\eta}$ and $C_{\eta}$ are numerical constant depending on $\eta$.
 Proof.
-Note that the expression of $f_{l}$ can be transformed as
-$f_{l}=\mathbf{e}_{l}^{*}\mathbf{Z}_{S^{0}}\mathbf{e}^{0},$
-where $\mathbf{e}_{l}\in{\mathbb{C}}^{n}$ is the vector that sets its $l$-th entry as 1 while others as 0, $S^{0}$ is a fixed set derived by Step 1 of gESP and $\mathbf{e}^{0}$ is the principle eigenvector of $\mathbf{Z}_{S^{0}}$.
-Employing Lemma 4 with fixed $S=S^{0}$, fixed $T=\{\mathbf{e}_{1},\mathbf{e}_{2},\cdots,\mathbf{e}_{n},\mathbf{e}^{0}\}$ and $g(k,s(p))=\frac{2}{\gamma}\sqrt{ks(p)}$, we derive that
-$\max_{l\in[n]}\left\lvert f_{l}-\mathbb{E}(\tilde{f}_{l})\right\rvert<\eta\times\frac{\gamma}{2\sqrt{ks(p)}}$
-holds true with probability exceeding $1-2\exp(-c_{\eta}m)$ with the sample complexity satisfying $m\geq C_{\eta}\frac{ks(p)}{\gamma^{2}}\log n$.
+The goal is to bound the element-wise error $\max_{l}|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$.
+We rely on the independence between the measurement vectors $\mathbf{a}_{i}$ in the current stage and the vector $\mathbf{e}^{0}$.
+We decompose the error into two terms as follows:
+$|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ $\leq\underbrace{|f_{l}-\tilde{f}_{l}|}_{\text{bias term}}+\underbrace{|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|}_{\text{fluctuation term}}$
+$=|\mathbf{e}_{l}^{*}(\mathbf{Z}-\tilde{\mathbf{Z}})\mathbf{e}^{0}|+|\mathbf{e}_{l}^{*}(\tilde{\mathbf{Z}}-\mathbb{E}[\tilde{\mathbf{Z}}])\mathbf{e}^{0}|.$ (A.17)
+On one hand, the bias term is controlled by the operator norm of the difference matrix $\Delta\mathbf{Z}=\mathbf{Z}-\tilde{\mathbf{Z}}$.
+Note that $\mathbf{e}^{0}$ is supported on $S^{0}$, and denote $U=S^{0}\cup\{l\}$ of size $|U|=p+1\leq 2k$.
+Then we have
+$|\mathbf{e}_{l}^{*}(\mathbf{Z}-\tilde{\mathbf{Z}})\mathbf{e}^{0}|$ $\makebox[61.83356pt][c]{$\leq$}\|(\mathbf{Z}-\tilde{\mathbf{Z}})_{U}\|\|\mathbf{e}_{l}\|_{2}\|\mathbf{e}^{0}\|_{2}$ (A.20)
+$\makebox[61.83356pt][c]{$\leq$}\|(\mathbf{Z}-\tilde{\mathbf{Z}})_{U}\|$ (A.23)
+$\makebox[61.83356pt][c]{$\overset{(a)}{\leq}$}C\sqrt{\frac{\log n}{m}},$ (A.26)
+where (a) follows from Lemma 5.
+On the other hand, conditioned on $S^{0}$ (and thus $\mathbf{e}^{0}$), we analyze the fluctuation term $|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|$.
+Recall that $\tilde{\mathbf{f}}=\tilde{\mathbf{Z}}\mathbf{e}^{0}$.
+We can expand the $l$-th entry as a sum of $m$ independent random variables:
+$\tilde{f}_{l}=\mathbf{e}_{l}^{*}\tilde{\mathbf{Z}}\mathbf{e}^{0}=\frac{1}{m}\sum_{i=1}^{m}\kappa_{i}(\mathbf{a}_{i}^{*}\mathbf{e}_{l})^{*}(\mathbf{a}_{i}^{*}\mathbf{e}^{0}),$ (A.27)
+where $\kappa_{i}=\frac{1}{2}-\exp(-y_{i}^{2}/\|\mathbf{x}\|^{2})$.
+Since we assume independence between the current samples $\{\mathbf{a}_{i}\}$ and the initialization $\mathbf{e}^{0}$ (data splitting), the terms in the summation are independent conditioned on $\mathbf{e}^{0}$.
+Let $u_{i}=\mathbf{a}_{i}^{*}\mathbf{e}_{l}$ and $v_{i}=\mathbf{a}_{i}^{*}\mathbf{e}^{0}$.
+Both $u_{i}$ and $v_{i}$ are standard complex Gaussian variables (potentially correlated if $l\in\text{supp}(\mathbf{e}^{0})$).
+The $i$-th summand is $X_{i}=\kappa_{i}\bar{u}_{i}v_{i}$.
+Note that the weights are bounded as $|\kappa_{i}|\leq 1/2$.
+The product of two Gaussian variables $\bar{u}_{i}v_{i}$ follows a sub-exponential distribution with bounded Orlicz norm $\|\bar{u}_{i}v_{i}\|_{\psi_{1}}\leq C\|u_{i}\|_{\psi_{2}}\|v_{i}\|_{\psi_{2}}\leq C^{\prime}$.
+Consequently, the weighted variable $X_{i}$ is sub-exponential with
+$\|X_{i}\|_{\psi_{1}}\leq|\kappa_{i}|\|\bar{u}_{i}v_{i}\|_{\psi_{1}}\leq C^{\prime\prime}.$ (A.28)
+Applying the scalar Bernstein inequality [23, Theorem 2.8.1] directly, we have
+$\mathbb{P}(|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|\leq t)\geq 1-2\exp\left(-cm\min\{t,t^{2}\}\right).$
+Setting $t=\sqrt{\frac{\log n}{m}}$ and taking the union bound over all $l\in[n]$, we get
+$\max_{l}|\tilde{f}_{l}-\mathbb{E}[\tilde{f}_{l}]|\leq C^{\prime}\sqrt{\frac{\log n}{m}}.$ (A.29)
+Combining (A.26) and (A.29), we have
+$|f_{l}-\mathbb{E}[\tilde{f}_{l}]|\leq C\sqrt{\frac{\log n}{m}}.$ (A.30)
+Therefore, when $m\geq C\eta^{-2}\gamma^{-2}ks(p)\log n$, the desired result holds with probability exceeding $1-n^{-c_{\eta}}$. ∎
+A-C Proof for Example 1
+For sufficiently large $k$, we have $k^{-\delta}\gg k^{-1/2}$ (since $\delta<1/2$), so the ordering in (49) is consistent with $x_{(1)}\geq x_{(2)}\geq\cdots\geq x_{(k)}$.
+First, we have
+$\|\mathbf{x}\|^{2}=k^{-2\delta}+(m-1)\cdot\frac{1}{4}k^{-2\delta}+(k-m)\cdot k^{-1}=\Theta(1),$ (A.31)
+because $(m-1)k^{-2\delta}=\Theta\!\big(k^{1/2-2\delta}\big)=o(1)$ for $\delta>1/4$, and $(k-m)k^{-1}=1-\Theta(k^{-1/2})$.
+We now compare the orders of magnitude for the quantities of interest:
+  • For $p=1$,
+    $s(1)=\frac{\|\mathbf{x}\|^{2}}{|x_{(1)}|^{2}}=\Theta(1)\cdot k^{2\delta}=\Theta\!\big(k^{2\delta}\big).$ (A.32)
+  • For $m=\lceil\sqrt{k}\rceil$,
+    $\sum_{j=1}^{m}|x_{(j)}|^{2}=k^{-2\delta}+(m-1)\cdot\frac{1}{4}k^{-2\delta}=\Theta\!\big(k^{1/2-2\delta}\big).$ (A.33)
+    Thus we have
+    $s^{2}(m)=\Theta\!\big(k^{4\delta-1}\big).$ (A.34)
+Consequently, for all $\delta\in(1/4,1/2)$, we have
+$\frac{s^{2}(m)}{s(1)}=\Theta\!\big(k^{2\delta-1}\big)\to 0,$ (A.35)
+and hence $\Theta(s^{2}(\lceil\sqrt{k}\rceil))<\Theta(s(1))$.
 Appendix B Proof for Proposition 1
@@
 We establish the upper and lower bounds for $\frac{1}{p}\sum_{j\in S^{0}}Z_{j}$ and $\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}$, respectively.
-$\frac{1}{p}\sum_{j\in S^{0}}Z_{j}\leq$ $~\frac{1}{p}\sum_{j\in S^{0}}\mathbb{E}\left[\tilde{Z}_{j}\right]+\left\lvert\frac{1}{p}\sum_{j\in S^{0}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
-$\leq$ $~\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{4p\|\mathbf{x}\|^{2}}+\frac{1}{16ps(p)},$ (B.1)
-$\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}\geq$ $~\frac{1}{p}\sum_{j\in\overline{S}}\mathbb{E}\left[\tilde{Z}_{j}\right]-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
-$=$ $\frac{\|\mathbf{x}_{\overline{S}}\|^{2}}{4p\|\mathbf{x}\|^{2}}-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathrm{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
+Applying Lemma 6 yields
+$\frac{1}{p}\sum_{j\in S^{0}}Z_{j}\leq$ $~\frac{1}{p}\sum_{j\in S^{0}}\mathbb{E}\left[\tilde{Z}_{j}\right]+\left\lvert\frac{1}{p}\sum_{j\in S^{0}}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
+$\leq$ $~\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{4p\|\mathbf{x}\|^{2}}+\frac{1}{16ps(p)}$ (B.1)
+and
+$\frac{1}{p}\sum_{j\in\overline{S}}Z_{j}\geq$ $~\frac{1}{p}\sum_{j\in\overline{S}}\mathbb{E}\left[\tilde{Z}_{j}\right]-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
+$=$ $\frac{\|\mathbf{x}_{\overline{S}}\|^{2}}{4p\|\mathbf{x}\|^{2}}-\left\lvert\frac{1}{p}\sum_{j\in\overline{S}}\left(Z_{j}-\mathbb{E}\left[\tilde{Z}_{j}\right]\right)\right\rvert$
 $\geq$ $~\frac{1}{4ps(p)}-\frac{1}{16ps(p)}=\frac{3}{16ps(p)}.$ (B.2)
@@
 $\frac{\|\mathbf{x}_{S^{0}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}.$ (B.5)
-Thus we complete the proof.
+This completes the proof.
 Appendix C Proof for Proposition 2
-For any set $\mathcal{T}$} satisfying $|{\mathcal{T}}|=p^{\prime}\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$, suppose $\mathbf{z}$ as the unit eigenvector corresponding to the largest eigenvalue, denoted as $\tau$, of
+For any set $\mathcal{T}$} satisfying $|{\mathcal{T}}|=p^{\prime}\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$, let $\mathbf{z}$ be the unit eigenvector corresponding to the largest eigenvalue, denoted as $\tau$, of
 $\mathbf{Z}_{{\mathcal{T}}}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{1}{2}-\exp\left(-\frac{y_{i}^{2}}{\lambda^{2}}\right)\right)(\mathbf{a}_{i})_{{\mathcal{T}}}(\mathbf{a}_{i})^{*}_{{\mathcal{T}}}.$
@@
 $\overset{(a)}{\leq}$ $\left\lVert\mathbf{Z}_{{\mathcal{T}}}-\frac{\mathbf{x}_{{\mathcal{T}}}\mathbf{x}^{*}_{{\mathcal{T}}}}{4\|\mathbf{x}\|^{2}}\right\rVert\|\mathbf{z}\|^{2}\overset{(b)}{\leq}\frac{\eta}{s(p)},$ (C.6)
-where $(a)$ is based on the relationship between matrix norm and vector norm, $(b)$ employs (A.9) with $g(k,s(p))=s(p)$ and the condition $\|\mathbf{z}\|=1$.
-The above inequality holds with probability at least $1-2e^{-c_{\eta}m}$ when $m\geq C_{\eta}p^{\prime}s^{2}(p)\log n$.
+where $(a)$ is based on the relationship between matrix norm and vector norm,
+and $(b)$ employs Corollary 4.
+The above inequality holds with probability at least $1-n^{-c_{\eta}}$ when $m\geq C_{\eta}p^{\prime}s^{2}(p)\log n$ since $|{\mathcal{T}}|=p^{\prime}$.
 Therefore, through simple transformation, (C.6) can be expressed as
 $|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\geq 4(\tau-\frac{\eta}{s(p)})\|\mathbf{x}\|^{2}.$ (C.7)
-Moreover, since $\tau$ is the largest eigenvalue of $\mathbf{Z}_{{\mathcal{T}}}$, we can estimate it as
+Moreover, since $\tau$ is the largest eigenvalue of the Hermitian matrix $\mathbf{Z}_{{\mathcal{T}}}$, we can estimate it as
 […]
@@
 $\overset{(a)}{\geq}$ $~-\frac{\eta}{s(p)}+\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}},$ (C.16)
-where $(a)$ comes from the definition of matrix norm and Lemma 4 ((A.9)) with g(k,s(p))=s(p)g(k,s(p))=s(p) [^5] In fact, since the matrix $\mathbf{Z}_{TT}-\frac{\mathbf{x}_{{\mathcal{T}}}\mathbf{x}^{*}_{{\mathcal{T}}}}{4\|\mathbf{x}\|^{2}}$ are Hermitian matrix, all of its eigenvalues are real and non-negative, allowing the inequality hold true.
-.
-Taking (C.16) into (C.7) yields
+where $(a)$ comes from the definition of matrix norm and Corollary 4. Taking (C.16) into (C.7) yields
 $|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\geq$ $~4\left(\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}}-2\frac{\eta}{s(p)}\right)\|\mathbf{x}\|^{2}$
 $=$ $~\|\mathbf{x}_{{\mathcal{T}}}\|^{2}-8\frac{\eta}{s(p)}\|\mathbf{x}\|^{2}\overset{(a)}{\geq}(1-16\eta)\|\mathbf{x}_{{\mathcal{T}}}\|^{2},$ (C.17)
-where $(a)$ comes from the assumption $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$. Let $\mu_{1}=\sqrt{(1-16\eta)}$, we complete the proof.
+where $(a)$ comes from the assumption $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq\frac{1}{2s(p)}$. Let $\sqrt{1-16\eta}=1/2$, we complete the proof.
 Appendix D Proof for Proposition 3
@@
 […]
-$\overset{(a)}{\geq}$ $~\frac{1}{4\|\mathbf{x}\|^{2}}\times\frac{\gamma\|\mathbf{x}\|}{2\sqrt{k}}\times\mu_{1}\sqrt{\frac{1}{2s(p)}}\|\mathbf{x}\|-|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ (D.7)
-$\overset{(b)}{>}$ $~\frac{\gamma\mu_{1}}{8\sqrt{2}}\sqrt{\frac{1}{ks(p)}}-\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}}=\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.10)
-where $(a)$ is from the definition of $S_{\gamma}$ and Proposition 2 and $(b)$ employs Lemma 6 with $\eta=\frac{\mu_{1}}{8\sqrt{2}}$ and holds with probability exceeding $1-4e^{-c_{\gamma,\mu_{1}}m}$ when $m\geq C_{\mu_{1}}\frac{ks(p)}{\gamma^{2}}\log n$.
+$\overset{(a)}{\geq}$ $~\frac{1}{4\|\mathbf{x}\|^{2}}\times\frac{1}{2}\sqrt{\frac{1}{2s(p)}}\|\mathbf{x}\|\times\frac{\gamma\|\mathbf{x}\|}{2\sqrt{k}}-|f_{l}-\mathbb{E}[\tilde{f}_{l}]|$ (D.7)
+$\overset{(b)}{>}$ $~\frac{\gamma}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}}-\frac{\gamma}{32\sqrt{2}}\sqrt{\frac{1}{ks(p)}}=\frac{\gamma}{32\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.10)
+where $(a)$ is from the definition of $S_{\gamma}$ and Proposition 2, and $(b)$ employs Lemma 7 with $\eta=\frac{1}{16\sqrt{2}}$ and holds with probability exceeding $1-n^{-c}$ when $m\geq C\gamma^{-2}ks(p)\log n$.
 On the other hand, for $l\notin{\text{supp}(\mathbf{x})}$, we have
-$|f_{l}|\leq|\mathbb{E}[\tilde{f}_{l}]|+|f_{l}+\mathbb{E}[\tilde{f}_{l}]|\overset{(a)}{\leq}\frac{\gamma\mu_{1}}{16\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.11)
-where $(a)$ comes from (D.1) and Lemma 6 with $\eta=\frac{\mu_{1}}{8\sqrt{2}}$.
+$|f_{l}|\leq|\mathbb{E}[\tilde{f}_{l}]|+|f_{l}+\mathbb{E}[\tilde{f}_{l}]|\overset{(a)}{\leq}\frac{\gamma}{32\sqrt{2}}\sqrt{\frac{1}{ks(p)}},$ (D.11)
+where $(a)$ comes from (D.1) and Lemma 7 with $\eta=\frac{1}{16\sqrt{2}}$.
 By combining (D.10) and (D.11), we claim that all indices in $S_{\gamma}$ are selected in this case.
 Therefore, we can derive that
-$\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$
+$\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\sum_{i\notin S_{\gamma}}\frac{|x_{i}|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}.$ (D.12)
 Appendix E Proof of Proposition 4
-Let $\mathbf{z}$ with $\|\mathbf{z}\|=\lambda$ be the eigenvector corresponding to the largest eigenvalue of $\mathbf{Z}_{{\mathcal{T}}}$.
-Following the same analysis as in Appendix C with the set $\mathcal{T}$} satisfying $|{\mathcal{T}}|\leq k$ and $\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}\geq 1-\frac{\gamma^{2}}{4}$, we obtain similar conclusion to (C.7) and (C.16) as
-$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}$ $\geq 4(\tau-\eta_{2})\|\mathbf{z}\|^{2}\|\mathbf{x}\|^{2},$ (E.1)
-$\tau$ $\geq\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{4\|\mathbf{x}\|^{2}}-\eta_{2},$ (E.2)
-where we employ $\|\mathbf{z}\|=\lambda$ and Lemma 4 with $g(k,s(p))=1$.
-Hence (E.1) and (E.2) hold with probability at least $1-2e^{-c_{\gamma,\eta_{2}}m}$ if $m\geq C_{\gamma,\eta_{2}}k\log n$.
-Suppose $\kappa_{0}=\frac{\|\mathbf{x}_{{\mathcal{T}}}\|^{2}}{\|\mathbf{x}\|^{2}}$, taking (E.2) into (E.1) yields
-$|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|^{2}\overset{(a)}{\geq}\left(\kappa_{0}-8\eta_{2}\right)(1-\eta_{1})\|\mathbf{x}\|^{4}\overset{\triangle}{=}\mu_{2}^{2}\|\mathbf{x}\|^{4},$ (E.3)
-where $(a)$ employs Lemma 3 with $\kappa=1$ and $\beta=\eta_{1}$.
-In the following, We derive the estimation for $\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})$.
-From the property of inner product, we have
-$\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}=$ $~\|\mathbf{x}_{{\mathcal{T}}}\|^{2}+\|\mathbf{z}\|^{2}-2|\mathbf{x}^{*}_{{\mathcal{T}}}\mathbf{z}|$
-$\leq$ $~(1+\eta_{1}+\kappa_{0})\|\mathbf{x}\|^{2}-2\mu_{2}\|\mathbf{x}\|^{2}.$ (E.4)
-Take $\eta_{1}=1-\kappa_{0}$ and $8\eta_{2}=\kappa_{0}-\kappa_{0}^{2}$, (E.4) becomes
-$\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}\leq(2-2\sqrt{\kappa_{0}^{3}})\|\mathbf{x}\|^{2}.$ (E.5)
-Therefore, for any $0<\delta_{0}<1$, take $\gamma$ as
-$1-\frac{\gamma^{2}}{4}=\sqrt[3]{\left(\frac{2-\delta_{0}^{2}}{2}\right)^{2}},$
-and the result (E.5) becomes
-$\text{dist}(\mathbf{x}_{{\mathcal{T}}},z)\leq\delta_{0}\|\mathbf{x}\|.$ (E.6)
-Finally, we have
-$\text{dist}(\mathbf{x},\mathbf{z})^{2}$~ $\makebox[29.6192pt][c]{$=$}~\text{dist}(\mathbf{x}_{{\mathcal{T}}},\mathbf{z})^{2}+\|\mathbf{x}^{\perp}_{{\mathcal{T}}}\|^{2}$
-$\makebox[29.6192pt][c]{$\overset{(a)}{\leq}$}~\delta_{0}^{2}\|\mathbf{x}\|^{2}+\frac{\gamma^{2}}{4}\|\mathbf{x}\|^{2}=\left(\delta_{0}^{2}+\frac{\gamma^{2}}{4}\right)\|\mathbf{x}\|^{2},$
-where $(a)$ comes from (E.6) and the definition of $\mathcal{T}$}.
-Moreover, condition (51) indicates that $\delta\geq\sqrt{\delta_{0}^{2}+\frac{\gamma^{2}}{4}}\in[0,1]$.
-Therefore, we complete the proof of Proposition 4.
+Recall that in Step 4 of the algorithm, $\mathbf{z}$ is defined as the eigenvector of $\mathbf{Z}_{S^{1}}$ corresponding to the largest eigenvalue, scaled such that $\|\mathbf{z}\|_{2}=\lambda$, where $\lambda=\sqrt{\frac{1}{m}\sum_{i=1}^{m}y_{i}^{2}}$.
+Let $\hat{\mathbf{u}}=\mathbf{z}/\|\mathbf{z}\|_{2}$ be the estimated direction, and $\mathbf{u}^{*}=\mathbf{x}_{S^{1}}/\|\mathbf{x}_{S^{1}}\|_{2}$ be the true direction on the subset $S^{1}$.
+By the definition of the distance, we have
+$\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})$ $=\min_{\phi\in[0,2\pi)}\|\mathbf{z}-e^{j\phi}\mathbf{x}_{S^{1}}\|_{2}$
+$\leq\|\mathbf{z}-e^{j\phi^{*}}\mathbf{x}_{S^{1}}\|_{2}$
+$=\Big\|\lambda\hat{\mathbf{u}}-e^{j\phi^{*}}\|\mathbf{x}_{S^{1}}\|_{2}\mathbf{u}^{*}\Big\|_{2},$ (E.1)
+where $\phi^{*}$ is the optimal phase aligning $\hat{\mathbf{u}}$ and $\mathbf{u}^{*}$.
+By adding and subtracting $\lambda e^{j\phi^{*}}\mathbf{u}^{*}$, we can decompose the error into a norm component and a directional component via the triangle inequality:
+$~~~~\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})$
+$\leq\Big\|\lambda\hat{\mathbf{u}}-\lambda e^{j\phi^{*}}\mathbf{u}^{*}\Big\|_{2}+\Big\|\lambda e^{j\phi^{*}}\mathbf{u}^{*}-\|\mathbf{x}_{S^{1}}\|_{2}e^{j\phi^{*}}\mathbf{u}^{*}\Big\|_{2}$
+$=\lambda\underbrace{\|\hat{\mathbf{u}}-e^{j\phi^{*}}\mathbf{u}^{*}\|_{2}}_{\text{directional error}}+\underbrace{|\lambda-\|\mathbf{x}_{S^{1}}\|_{2}|}_{\text{norm error}}.$ (E.2)
+On one hand, for the directional error, the expected matrix is $\mathbb{E}[\tilde{\mathbf{Z}}_{S^{1}}]=\frac{\mathbf{x}_{S^{1}}\mathbf{x}_{S^{1}}^{*}}{4\|\mathbf{x}\|^{2}}$.
+The eigengap between the largest eigenvalue and the second largest (which is 0) is $\delta_{gap}=\frac{\|\mathbf{x}_{S^{1}}\|^{2}}{4\|\mathbf{x}\|^{2}}$.
+According to the Davis-Kahan $\sin\Theta$ theorem [7], the distance between the eigenvectors is bounded by the ratio of the perturbation to the eigengap:
+$\|\hat{\mathbf{u}}-e^{j\phi^{*}}\mathbf{u}^{*}\|_{2}$ $\leq\frac{\sqrt{2}\|\mathbf{Z}_{S^{1}}-\mathbb{E}[\tilde{\mathbf{Z}}_{S^{1}}]\|}{\|\mathbf{x}_{S^{1}}\|^{2}/(4\|\mathbf{x}\|^{2})}$
+$\overset{(a)}{\leq}\frac{4\sqrt{2}\|\mathbf{x}\|^{2}}{\|\mathbf{x}_{S^{1}}\|^{2}}C\sqrt{\frac{k\log n}{m}},$ (E.3)
+where $(a)$ comes from Corollary 4, with probability at least $1-2e^{-ck\log n}$.
+On the other hand, for the norm error, we need to bound $|\lambda-\|\mathbf{x}_{S^{1}}\|_{2}|$. By the triangle inequality, we have
+$|\lambda-\|\mathbf{x}_{S^{1}}\|_{2}|\leq|\lambda-\|\mathbf{x}\|_{2}|+|\|\mathbf{x}\|_{2}-\|\mathbf{x}_{S^{1}}\|_{2}|.$ (E.4)
+First, from Lemma 2, we know that $|\lambda^{2}-\|\mathbf{x}\|^{2}|\leq C\sqrt{\frac{\log n}{m}}\|\mathbf{x}\|^{2}$. Using the inequality $|\sqrt{a}-\sqrt{b}|\leq\frac{|a-b|}{\sqrt{a}}$, we have
+$|\lambda-\|\mathbf{x}\|_{2}|\leq\frac{|\lambda^{2}-\|\mathbf{x}\|^{2}|}{\|\mathbf{x}\|_{2}}\leq C\sqrt{\frac{\log n}{m}}\|\mathbf{x}\|_{2}.$ (E.5)
+When $m>c\log n$, it holds that
+$\lambda\leq\left(1+C\sqrt{\frac{\log n}{m}}\right)\|\mathbf{x}\|_{2}\leq 2\|\mathbf{x}\|_{2}.$ (E.6)
+Second, from Proposition 3, we know $S^{1}$ contains the dominant entries $S_{\gamma}$, implying
+$\|\mathbf{x}_{S^{1}}\|_{2}\geq\sqrt{1-\gamma^{2}/4}\|\mathbf{x}\|_{2}.$ (E.7)
+We can always choose small $\gamma\leq 1/2$ such that $\sqrt{1-\gamma^{2}/4}\geq 1-\gamma/2$. Thus, we have
+$|\|\mathbf{x}\|_{2}-\|\mathbf{x}_{S^{1}}\|_{2}|\leq\frac{\gamma}{2}\|\mathbf{x}\|_{2}.$ (E.8)
+Collecting the above bounds (E.3) – (E.8) with constants absorbed into $C_{1}$, $C_{2}$, we have
+$~~~~\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})$ $\leq C_{1}\left(1-\frac{\gamma^{2}}{4}\right)^{-1}\|\mathbf{x}\|_{2}\sqrt{\frac{k\log n}{m}}$
+$+$~ $\left(C_{2}\sqrt{\frac{\log n}{m}}+\frac{\gamma}{2}\right)\|\mathbf{x}\|_{2}.$ (E.9)
+Choose a small constant $\gamma$ such that $\left(1-\frac{\gamma^{2}}{4}\right)^{-1}\leq 2$. We have
+$\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})\leq C\sqrt{\frac{k\log n}{m}}\|\mathbf{x}\|_{2}+\frac{\gamma}{2}\|\mathbf{x}\|_{2}.$ (E.10)
+Finally, considering the total distance to the original signal $\mathbf{x}$:
+$\text{dist}(\mathbf{z},\mathbf{x})$ $\leq\text{dist}(\mathbf{z},\mathbf{x}_{S^{1}})+\|\mathbf{x}_{S^{1}}-\mathbf{x}\|_{2}$
+$\leq C\|\mathbf{x}\|\sqrt{\frac{k\log n}{m}}+\gamma\|\mathbf{x}\|.$ (E.11)
+To ensure $\text{dist}(\mathbf{z},\mathbf{x})\leq\delta\|\mathbf{x}\|$, we set $\gamma<\frac{1}{2}\delta$.
+Proposition 3 implies a scaling dependency of $m\propto\gamma^{-2}$.
+Consequently, setting $\gamma<\frac{1}{2}\delta$ requires $m\propto\delta^{-2}$.
+Crucially, this sample complexity is sufficient to bound the first term in (E.11) by $\frac{1}{2}\delta\|\mathbf{x}\|$ as well.
+Thus, the condition is satisfied provided that
+$m\geq C\delta^{-2}k\log n.$ (E.12)